### 17 Apr 2026 v2

# A geometric approach to exponentially small splitting: The generic zero-Hopf bifurcation of co-dimension two

Abstract.
In this paper, we consider the unfolding of the real-analytic and generic zero-Hopf bifurcation of co-dimension two. It is well-known that in an open set of parameter space the splitting of one-dimensional stable and unstable manifolds is beyond all orders. This paper provides a new geometric dynamical-systems-oriented proof for the exponentially small splitting. As a novel aspect, we relate the exponentially small splitting to the lack of analyticity of center-like manifolds of generalized saddle-nodes. Moreover, the blowup method plays an important technical role as a systematic way to relate dynamics on different orders of magnitude. Finally, as our approach takes place in the (complexified) phase space, we do not rely on an explicit time-parametrization of the unperturbed heteroclinic connection. We therefore believe that our approach has general interest.
Keywords. Zero-Hopf bifurcation, blowup, exponentially small splitting, GSPT, generalized saddle-nodes.
Mathematics Subject Classification. 34C23, 34D15, 34C45, 37G10
1. Introduction
In this paper, we revisit the exponentially small splitting in the unfolding of the real-analytic zero-Hopf bifurcation. We recall that the zero-Hopf bifurcation [25] is defined for a vector-field $\Xi_{0}$ in $\mathbb{R}^{3}$ as a singularity with eigenvalues 0, $\pm i\omega$, $\omega\neq 0$, of the linearization. We will consider the generic case where the zero-Hopf bifurcation is co-dimension two [9], using $(\mu,\nu)\in\mathbb{R}^{2}$, $(\mu,\nu)\sim(0,0)$, to denote the unfolding parameters. In particular, by [1, Eq. (6)] we have the following normal form for the unfolding $\Xi_{\mu,\nu}$ of $\Xi_{0}$:
$x^{\prime}$ $=x^{2}-\mu+a(y^{2}+z^{2})+F(x,y,z,\mu,\nu),$ (1.1)
$y^{\prime}$ $=(\nu-bx)y+z+G(x,y,z,\mu,\nu),$
$z^{\prime}$ $=(\nu-bx)z-y+H(x,y,z,\mu,\nu),$
with $(\cdot)^{\prime}=\frac{d}{d\tau}$. Here $F,G$ and $H$ are real-analytic functions, defined in a neighborhood of $(x,y,z,\mu,\nu)=(0,0,0,0,0)$ in $\mathbb{C}^{5}$, and each function is third order:
$W(x,y,z,\mu,\nu)=\mathcal{O}(|(x,y,z,\mu,\nu)|^{3}),\quad W=F,G,H,$ (1.2)
with respect to $(x,y,z,\mu,\nu)\to(0,0,0,0,0)$. For the derivation of (1.1), we have used a re-parametrization of time that brings $\omega=1$, see Appendix A for details. Here we also show that it is without loss of generality to assume that
$F(x,0,0,\mu,\nu)=F_{0}x^{3}+\mathcal{O}(|(x,\mu,\nu)|^{4}),\quad F_{0}\in\mathbb{R},$ (1.3)
with respect to $(x,\mu,\nu)\to(0,0,0)$. We will as in [1] suppose that
$b>0.$
Then we have the following: Fix any $\delta\in(0,1)$ and let $\mathcal{W}$ denote the region of parameter space $(\mu,\nu)\in\mathbb{R}^{2}$, defined by
$\begin{cases}\mu=\epsilon^{2},\\ \nu=\epsilon b\sigma,\end{cases}\quad\sigma\in[-1+\delta,1-\delta],\quad 0<\epsilon<\epsilon_{0},$ (1.4)
with $0<\epsilon_{0}=\epsilon_{0}(\delta)\ll 1$, see Fig. 1. Then for any $(\mu,\nu)\in\mathcal{W}$, the system (1.1) has two saddle-foci $E^{\pm}(\mu,\nu)$ with $x$-components given by the form $\pm\epsilon(1+\mathcal{O}(\epsilon))$, respectively. For further details, see [1] or Section 2 below. There is a one-dimensional heteroclinic connection for $\epsilon=0$, as well as a formal one in powers of $\epsilon$. The splitting of the one-dimensional unstable and stable manifolds of $E^{\pm}$ is therefore a beyond-all-orders phenomena with respect to $\epsilon\rightarrow 0$. We are as in [1] concerned with an expression for their splitting for all $0<\epsilon\ll 1$.
Figure 1. Illustration of the region $\mathcal{W}\subset\mathbb{R}^{2}$ in the parameter $(\mu,\nu)$-plane. For any $(\mu,\nu)\in\mathcal{W}$, the system (1.1) has two saddle-foci $E^{\pm}(\mu,\nu)$. The splitting of the associated unstable and stable manifolds is beyond all orders as $\epsilon\to 0$ (with $\sigma\in(-1,1)$ fixed), see (1.4).
Remark 1.1.
For $a>0$, there is a related splitting problem associated with the remaining two-dimensional invariant manifolds (carrying focus-type dynamics) of $E^{\pm}(\mu,\nu)$, which is important for the existence of Shilnikov bifurcations, see [5] as well as [9] for Shilnikov bifurcations in the $C^{\infty}$-context. However, in this paper, we will only focus on the one-dimensional splitting problem, and will therefore not make any assumptions on the parameter $a$. For details on $(\mu,\nu,b)\notin\mathcal{W}\times\mathbb{R}_{+}$, we refer to [9, 25].
We now describe the result of [1] in further details. For this, consider (1.1) with $b>0$ and $(\mu,\nu)\in\mathcal{W}$ given by (1.4):
$x^{\prime}$ $=x^{2}-\epsilon^{2}+a(y^{2}+z^{2})+F(x,y,z,\epsilon),$ (1.5)
$y^{\prime}$ $=b(\sigma\epsilon-x)y+z+G(x,y,z,\epsilon),$
$z^{\prime}$ $=b(\sigma\epsilon-x)z-y+H(x,y,z,\epsilon),$
where we from now on use $W(x,y,z,\epsilon)$ to denote $W(x,y,z,\epsilon^{2},\epsilon b\sigma)$ for $W=F,G,H$. By assumption, we have
$W(x,y,z,\epsilon)=\mathcal{O}(|(x,y,z,\epsilon)|^{3}),\quad W=F,G,H.$ (1.6)
Now, with $b>0$ and $\sigma\in(-1,1)$ fixed, the main result of [1] can then be summarized as an asymptotic expression for the distance $d(\epsilon)$, $0<\epsilon\ll 1$, between the one-dimensional unstable and stable manifolds of $E^{\pm}(\epsilon):=E^{\pm}(\epsilon^{2},\epsilon b\sigma)$, respectively, at the fixed section $\{x=0\}$:
$d(\epsilon)=\epsilon^{-b}\mathrm{e}^{-\frac{\pi}{2}\epsilon^{-1}}\left(C_{0}+\mathcal{O}(1/\log\epsilon^{-1})\right),$ (1.7)
as $\epsilon\to 0$, for some constant $C_{0}\geq 0$ independent of $\epsilon$ and $\sigma$. Notice when comparing with [1] that our normal form corresponds to $\alpha_{0}=1$, $\alpha_{1}=\alpha_{3}=0$, $h_{0}=F_{0}$ and that our $C_{0}$ is therefore related to their $C^{*}$ by
$C_{0}={\mathrm{e}}^{\frac{\pi}{2}F_{0}}C^{*}.$
Generically, $C_{0}\neq 0$ and (1.7) therefore determines an asymptotic formula for the leading order splitting.
The proof of the expression (1.7) in [1] follows a well-known approach (initiated by Lazutkin for maps, see [35]): First, the authors scale the variables. In the present paper, we will use the following scaling
$\begin{cases}x=\epsilon x_{2},\\ y=\epsilon^{2}y_{2},\\ z=\epsilon^{2}z_{2},\end{cases}$ (1.8)
rather than the scaling used in [1]
$\begin{cases}x=\epsilon x_{2},\\ y=\epsilon y_{2},\\ z=\epsilon z_{2},\end{cases}$
but this difference is not essential. The transformation defined by (1.8) brings (1.5) into the following slow-fast form
$\dot{x}_{2}$ $=x_{2}^{2}-1+\epsilon^{2}a(y_{2}^{2}+z_{2}^{2})+\epsilon F_{2}(x_{2},\epsilon y_{2},\epsilon z_{2},\epsilon),$ (1.9)
$\epsilon\dot{y}_{2}$ $=z_{2}+\epsilon b(-x_{2}+\sigma)y_{2}+\epsilon G_{2}(x_{2},\epsilon y_{2},\epsilon z_{2},\epsilon),$
$\epsilon\dot{z}_{2}$ $=-y_{2}+\epsilon b(-x_{2}+\sigma)z_{2}+\epsilon H_{2}(x_{2},\epsilon y_{2},\epsilon z_{2},\epsilon),$
where $\frac{d}{dt}=\dot{(\cdot)}=\epsilon^{-1}(\cdot)^{\prime}$ and
$W_{2}(x_{2},\epsilon y_{2},\epsilon z_{2},\epsilon):=\epsilon^{-3}W(\epsilon x_{2},\epsilon(\epsilon y_{2}),\epsilon(\epsilon z_{2}),\epsilon),\quad W=F,G,H.$
Here $F_{2},G_{2},H_{2}$ all have real-analytic extensions to $\epsilon=0$ by (1.6). For $\epsilon=0$, we obtain $\dot{x}_{2}=x_{2}^{2}-1$, $y_{2}=z_{2}=0$, having a heteroclinic connection given by
$x_{2}=-\tanh(t)\to\mp 1\mbox{ for $t\to\pm\infty$, respectively}.$ (1.10)
The idea of [1] is then to parameterize the unstable and stable manifolds of $E^{\pm}(\epsilon)$ in terms of $t\in\mathbb{C}$ up close to the poles of (1.10) at $t=\pm i\frac{\pi}{2}$. In particular, upon setting
$t=\pm i\frac{\pi}{2}+\epsilon t_{2}$ (1.11)
with $t_{2}=\mathcal{O}(1)$, a so-called inner problem is defined for $\epsilon\to 0$, and it is shown that the unstable and stable manifolds lie close to special invariant manifold solutions of this inner problem. In this way, given that the distance between the invariant manifolds is known near the poles (with $t_{2}$ fixed, see (1.11)), the desired separation of the unstable and stable manifolds at $t=0$ (corresponding to $x=0$ by (1.10)) is obtained through the solution of a certain linear boundary value problem (obviously up to many technical details; see also [24, pp. 673–674] for a related summary of the approach).
The constant $C_{0}$ in (1.7) (known as a Stokes constant) is essentially a measure of the separation of the invariant manifold solutions of the inner problem.
The approach in [1], which is more functional-analytic in nature than dynamical-systems-oriented, has been successful in many important problems with exponentially small splitting, see e.g. [2, 3, 4, 6, 21, 24] and references therein. However, the point of departure (also within the context of maps [22, 35]) is always a parametrization of the unperturbed solution and its singularities in the complex plane, which is not only a strong assumption in a general context, but also an unusual one from the perspective of dynamical systems theory.
1.1. Main results
In this paper, we provide an alternative method for the asymptotic splitting of the one-dimensional invariant manifolds. In particular, our approach does not use complex time and the explicit solution of the $\epsilon=0$ limit, but instead we work exclusively in the complexified phase space $(x,y,z)\in\mathbb{C}^{3}$, parametrizing our invariant manifolds as graphs over $x$. Our approach is therefore more in tune with dynamical systems theory. It is our anticipation that the method will be useful in problems where explicit time parametrizations and their singularities are unknown.
[…]
We summarize our main results as follows (for further details, we refer to the main text below):
Theorem 1.2.
Consider (1.5), with $F$, $G$ and $H$ real-analytic on a neighborhood of $(x,y,z,\epsilon)=(0,0,0,0)$, satisfying (1.2), (1.3), and
$\mbox{$b>0$ \, and \, $-1<\sigma<1$},$ (1.12)
[…]
Let $m^{\pm}(0,\epsilon)\in\mathbb{R}^{2}$ denote the $(y,z)$-components of the first intersection of $\mathbf{W}_{\mathrm{loc}}^{u}(E^{+}(\epsilon))$ and $\mathbf{W}_{\mathrm{loc}}^{s}(E^{-}(\epsilon))$ with $\{x=0\}$, respectively. Then $m^{\pm}(0,\epsilon)$ are well-defined for all $\epsilon\in(0,\epsilon_{0})$ and the splitting is determined by
$m^{+}(0,\epsilon)-m^{-}(0,\epsilon)=\epsilon^{-b}{\mathrm{e}}^{-\frac{\pi}{2}\epsilon^{-1}}\operatorname{Re}\begin{pmatrix}{\mathrm{e}}^{iF_{0}\log\epsilon^{-1}+\phi_{0}(\epsilon)+\epsilon\phi_{1}(\epsilon)}\\ i{\mathrm{e}}^{iF_{0}\log\epsilon^{-1}+\phi_{0}(\epsilon)}\end{pmatrix}$, (1.13)
where $\phi_{\alpha}:[0,\epsilon_{0})\rightarrow\mathbb{C}$, $\alpha\in\{0,1\}$, are $C^{\infty}$-smooth functions. In particular, $m^{+}(0,\epsilon)\neq m^{-}(0,\epsilon)$ for all $\epsilon\in(0,\epsilon_{0})$.
Remark 1.3.
Through elementary calculations based on (1.13), we find that the distance $d(\epsilon):=|m^{+}(0,\epsilon)-m^{-}(0,\epsilon)|$ takes the following form
$d(\epsilon)=\epsilon^{-b}{\mathrm{e}}^{-\frac{\pi}{2}\epsilon^{-1}}\left(C_{0}+\mathcal{O}(\epsilon)\right),\quad C_{0}:={\mathrm{e}}^{\operatorname{Re}(\phi_{0}(0))}>0.$
Here $\mathcal{O}(\epsilon)$ is a $C^{\infty}$-smooth function of $(\cos(F_{0}\log\epsilon^{-1}),\sin(F_{0}\log\epsilon^{-1}),\epsilon)$.
Remark 1.4.
The unperturbed invariant manifolds $(y,z)=\psi_{0}^{\pm}(x)$, $x\in S^{\pm}$, are obtained by Borel-Laplace. In particular, the manifolds are analytic if and only if the Borel transform $\mathcal{B}(\psi_{0})(w)$ defines an entire function with at most exponential growth for $w\rightarrow\infty$, see [8, 16]. However, the linear part of the equation for $\mathcal{B}(\psi_{0})(w)$ has poles at $w=\pm i$ and this is also the general case for the actual Borel transform of $\psi_{0}^{\pm}$ (for the full nonlinear equation, see also [14, 31, 37, 43]).
Having said that, our approach does not require us to assume that the unperturbed manifolds are nonanalytic. In fact,
$m^{+}(0,\epsilon)-m^{-}(0,\epsilon)=\epsilon^{-b}{\mathrm{e}}^{-\frac{\pi}{2}\epsilon^{-1}}\operatorname{Re}\begin{pmatrix}\widetilde{\phi}_{0}(\epsilon){\mathrm{e}}^{iF_{0}\log\epsilon^{-1}+\epsilon\phi_{1}(\epsilon)}\\ i\widetilde{\phi}_{0}(\epsilon){\mathrm{e}}^{iF_{0}\log\epsilon^{-1}}\end{pmatrix}$, (1.14)
would be a more general statement for the splitting, with $\widetilde{\phi}_{0}(0)\neq 0$ if the unperturbed manifolds are nonanalytic at the origin. More generally, we have
$\begin{cases}\frac{\partial^{\alpha}\widetilde{\phi}_{0}}{\partial\epsilon^{\alpha}}(0)=0,\quad\forall\,\alpha\in\{0,1\ldots,\beta-1\},\\ \frac{\partial^{\beta}\widetilde{\phi}_{0}}{\partial\epsilon^{\beta}}(0)\neq 0,\end{cases}$
provided that $\beta\in\mathbb{N}$ is the smallest number for which the higher order “corrections” $\psi_{1,\beta}^{\pm}$, to the unperturbed manifold, are nonanalytic at the origin, see Lemma 4.3. These statements are simple consequences of (5.26) below. We therefore speculate whether the following holds true: Either
  1. (1)
    the formal series expansion of the connection converges and $d(\epsilon)=0$ for all $0<\epsilon\ll 1$,
or
  1. (2)
    there is some $\beta\in\mathbb{N}_{0}$ and some $C_{\beta}\neq 0$ such that
    $d(\epsilon)=\epsilon^{-b+\beta}{\mathrm{e}}^{-\frac{\pi}{2}\epsilon^{-1}}\left(C_{\beta}+\mathcal{O}(\epsilon)\right).$
This conjecture relates to the existence of “ultra-exponentially small splitting of separatrices” as discussed in [20]. We leave this open to future work.
1.2. A summary of our approach
A central technical aspect of the proof of Theorem 1.2, is that we view (1.8) as a local version of the blowup transformation
$r\geq 0,\,(\breve{x},\breve{y},\breve{z},\breve{\epsilon})\in\mathbb{S}^{3}\mapsto\begin{cases}x=r\breve{x},\\ y=r^{2}\breve{y},\\ z=r^{2}\breve{z},\\ \epsilon=r\breve{\epsilon}.\end{cases}$ (1.15)
(Here we refrain from using the standard notation (in blowup) $(\overline{x},\overline{y},\overline{z},\overline{\epsilon})\in\mathbb{S}^{3}$ for points on the sphere (see [33]), since we would like to reserve the bar for conjugation of complex numbers, and therefore use $(\breve{x},\breve{y},\breve{z},\breve{\epsilon})\in\mathbb{S}^{3}$ instead). In particular, we use the associated $\breve{x}=1$-chart, with chart-specific coordinates $(r_{1},y_{1},z_{1},\epsilon_{1})$ defined by
$\begin{cases}x=r_{1},\\ y=r_{1}^{2}y_{1},\\ z=r_{1}^{2}z_{1},\\ \epsilon=r_{1}\epsilon_{1},\end{cases}$ (1.16)
to extend the unstable and stable manifolds, which are naturally parameterized in compact subsets of the $(x_{2},y_{2},z_{2})$-space as graphs $(y_{2},z_{2})=m_{2}^{\pm}(x_{2},\epsilon)$ over $x_{2}\in\mathcal{X}_{2}^{\pm}$, see (1.9), to order $x=\mathcal{O}(1)\in\mathbb{C}$, $|x|>0$ sufficiently small. For the extension, we apply the flow in appropriate normal form coordinates, obtained by following the approach in [29], using two separate notions of formal invariant manifolds: (i) formal series with respect to $r_{1}$ with $\epsilon_{1}$-dependent coefficients, and (ii) formal series with respect to $\epsilon_{1}$ with $r_{1}$-dependent coefficients (with $r_{1}\in S^{\pm}\subset\mathbb{C}$). Our result shows that the unperturbed manifolds give the leading order approximation of the unstable and stable manifolds $(y,z)=m^{\pm}(x,\epsilon)$, for $x\in S^{\pm}$, $|x|>0$ small enough, respectively, as $\epsilon\to 0$. Therefore, whenever the unperturbed manifolds are distinct on the common domain $S^{+}\cap S^{-}$, we obtain an $\mathcal{O}(1)$-separation of the unstable and stable manifolds for $x\in S^{+}\cap S^{-}$, $|x|\geq\xi>0$ small enough, for $0<\epsilon\ll 1$. To determine the separation of the invariant manifolds at $\{x=0\}$, we then derive a linear equation for the difference $(\Delta y,\Delta z)=(m^{+}-m^{-})(x,\epsilon)$ (as in [1]) for $x$ on the imaginary axis $x\in i\mathbb{R}$, $|x|\geq 0$ small enough. As a novel aspect, we identify this set of equations as a slow-fast system with a normally hyperbolic critical manifold (of saddle-type) for $\epsilon\to 0$. Therefore, in contrast to the functional-analytic approach of [1], we use Geometric Singular Perturbation Theory (GSPT) [19, 28] to study this linear system and derive a Fenichel-type normal form [28, Section 5], bringing the linear system to diagonal form. In this way, the separation can easily be integrated from $x=\pm i\delta$, $\delta>0$ fixed small enough, to $x=0$. Here we also use that the equations are real-analytic: $(x,y,z)(t)$ is a solution of (1.5), $\epsilon>0$, if only if $(\overline{x},\overline{y},\overline{z})(t)$ is a solution, where the bar denotes the complex conjugate. We illustrate our approach in Fig. 2.
We believe that our approach to study the difference $(\Delta y,\Delta z)$ has general interest. Besides the zero-Hopf bifurcations of higher co-dimensions discussed below, we are presently pursuing the same approach to generalize the results in [26] to general delayed-Hopf phenomena (see [40, 41]).
Figure 2. Illustration of our approach. We blow up $(x,y,z,\epsilon)=(0,0,0,0)$ to a complex 33-sphere $\mathbb{S}^{3}$, see (1.15), which we here indicate in the $(\operatorname{Re}(x),\operatorname{Im}(x),\epsilon)$-projection. On top of the sphere ($\breve{\epsilon}>0$), we use the (scaled) coordinates $(x_{2},y_{2},z_{2})$, see (1.8), to describe the existence of the saddle-foci $E^{\pm}(\epsilon)$, $0<\epsilon\ll 1$, and their one-dimensional invariant manifolds as graphs $(y,z)=m^{\pm}(x,\epsilon)$ over $x_{2}=\epsilon^{-1}x\in\mathcal{X}^{\pm}\subset\mathbb{C}$, for all $0<\epsilon\ll 1$. The orange curves are orbits of $\dot{x}_{2}=x_{2}^{2}-1$, $x_{2}\in\mathbb{C}$, and $\mathcal{X}^{\mp}$ is any fixed compact set in the basin of attraction for $x_{2}=-1$ and $x_{2}=1$ (for the backward flow), respectively. A central objective of our approach is to extend the unstable and stable manifolds of $E^{\pm}(\epsilon)$, respectively, by working in the coordinates $(r_{1},y_{1},z_{1},\epsilon_{1})$, see (4.5), so that they can be compared with the unperturbed invariant manifolds that are defined as graphs over $x\in S^{\pm}$ (blue and green, respectively) within $\epsilon=0$. Our results show that if the unperturbed manifolds are non-analytic, then there is an $\mathcal{O}(1)$-splitting of the unstable and stable manifolds for $x\in S^{+}\cap S^{-}$, $|x|>0$ small enough, as $\epsilon\to 0$. Finally, we integrate the difference $(\Delta y,\Delta z)=(m^{+}-m^{-})(x,\epsilon)$ down along the imaginary axis $x\in i\mathbb{R}$. In this step, we view the equation for the difference as a slow-fast system with a normally hyperbolic critical manifold, which we then treat by GSPT (through a Fenichel-type normal form).
1.3. Discussion
An interesting aspect of our approach is that we do not use desingularization in association with our blowup. Essentially blowup is only used to compactify the $(x_{2},y_{2},z_{2})$-space. Hence we expect that our approach can also be applied to splitting problems associated with discrete time systems (where desingularization has been an obstruction in general, see [27]).
In [32], which can be seen as a follow-up paper, the present author studies an $\epsilon$-family of systems with degenerate zero-Hopf bifurcations for $\epsilon=0$ of arbitrary co-dimension $\kappa\in\mathbb{N}$, $\kappa\geq 3$, using the same approach. Here we find $2(\kappa-1)$ unperturbed invariant manifolds and the separation of these manifolds on common domains lead to exponentially small phenomena of $\kappa-1$ (formal) heteroclinic connections. In [32], we therefore generalize the last step of the method for dealing with the equation for the differences $(\Delta y,\Delta z)(x,\epsilon)$. For this, we use the notion of elliptic and hyperbolic paths in $x\in\mathbb{C}$ in a systematic way (these curves are also known as Stokes and anti-Stokes paths, respectively, see e.g. [41]). Such paths $x(s)\in\mathbb{C}$, $s\in\mathbb{R}$, in the complex plane are characterized by the linear part of the $(y,z)(s)$-subsystem being hyperbolic (expanding/contracting) and elliptic (oscillatory), respectively. In the present case of (1.5), these paths correspond to orbits of $\dot{x}_{2}=x_{2}^{2}-1$ and $\dot{x}_{2}=i(x_{2}^{2}-1)$, respectively. The results of [32] show that the $j$th splitting, $j\in\{1,\ldots,\kappa-1\}$, takes the form
$\epsilon^{e(\kappa)}{\mathrm{e}}^{-\epsilon^{1-\kappa}T_{j}}\left(C_{j}+\mathcal{O}(\epsilon)\right),\quad T_{j}>0,\,e(\kappa)\in\mathbb{R},$
under a nondegeneracy condition. Here $\{T_{j}\}_{j=1}^{\kappa-1}$ are finite blowup times for $(\kappa-1)$-many special unbounded solutions $\{\mathcal{H}_{j}\}_{j=1}^{\kappa-1}\subset\mathbb{C}$, respectively, of the reduced problem in imaginary time: $\dot{x}_{2}=iQ_{2}(x_{2})$ with $Q_{2}$ a polynomial of degree $\kappa$. All roots $\{q_{j}\}_{j=1}^{\kappa}$ of $Q_{2}$ are assumed to be real and simple. Then $T_{j}>0$ is expressed as
$T_{j}=\left|\sum_{l=1}^{j}\frac{\pi}{Q^{\prime}(q_{l})}\right|.$
We refer to [32] for further details. For $\kappa=2$ and $Q_{2}(x_{2})=x_{2}^{2}-1$, one has $\mathcal{H}_{1}=i\mathbb{R}$ and $T_{1}=\frac{\pi}{2}$ in agreement with (1.7). Now, the blowup in imaginary time is intrinsically related to singularities of time parametrizations of connecting orbits, see [20], but the time-parametrizations are not used explicitly in [32]; the construction is purely based upon phase space methods of dynamical systems theory.
For the generalization of our approach in [32] to higher co-dimension, we would like mention [26, 40, 41] as inspirations. In particular, the reference [26] used GSPT [19, 28] and blowup [18, 33] to describe exponentially small phenomena in simple model systems of the form $\epsilon z^{\prime}(t)=(t+i)z+\epsilon h(t,z,\epsilon)$ for slow passage through a Hopf bifurcation using hyperbolic paths. On the other hand, [40, 41] used elliptic paths to estimate the delay of general systems with slow passage through a Hopf bifurcation.
We emphasize that there is a different (but related) notion of Stokes and anti-Stokes lines of (exponential) asymptotic expansions of analytic functions, see [7, 10]. Recently, methods from exponential asymptotics have been used to describe exponentially small phenomena in nonlinear equations, see e.g. [10, 11, 36] and references therein. Although this approach appears to be more formal, it has been extremely successful. In fact, the results of [11], which show excellent agreement with numerical computations, presently seem to escape rigorous methods. The reference [10] has a summary of the approach: First, solutions are expanded in formal power series with respect $\epsilon$, having $t$-dependent coefficients $\phi_{\alpha}(t)$, $\alpha\in\mathbb{N}$. Next, one applies a WKB-type ansatz for the behaviour of $\phi_{\alpha}$ as $\alpha\to\infty$. This in turn induces an optimal truncation of the asymptotic series from which the dominant term of the exponentially small terms can be determined. A similar approach was used in the influential paper [34] by Kruskal and Segur on nonexistence of connecting orbits in a model for crystal growth.
Finally, we would like to mention [30] as an inspiration for this work. This paper studies weak analytic invariant manifolds of hyperbolic nodes close to saddle-nodes, which also relates to exponentially small phenomena. In summary, [30] shows (following [39]) that if the center manifold at the saddle-node is non-analytic, then the weak analytic manifolds can be controlled, see also [42]. The unperturbed invariant manifolds in the present paper are also center-like manifolds, being graphs over the zero-eigenspace. I therefore find that [30] bears some important similarities with the present paper.
1.4. Overview
The paper is organized as follows: In Section 2, we consider the $\epsilon=0$ limit and describe the unperturbed invariant manifolds, see Lemma 2.1. Subsequently, in Section 3 we consider (1.9) with $0<\epsilon\ll 1$ and describe the equilibria $E^{\pm}(\epsilon)$ and their one-dimensional invariant manifolds as graphs over large but fixed compact domains $x_{2}\in\mathcal{X}_{2}^{\pm}\subset\mathbb{C}$ (corresponding to $x=\mathcal{O}(\epsilon)$ by (1.8)) for $0<\epsilon\ll 1$, see Proposition 3.3. In Section 4, we describe our blowup approach. In particular, in this section we extend the unstable and stable manifolds to $x=\mathcal{O}(1)$, by working in the $\breve{x}=1$-chart associated with (1.15), see (1.16), and compare these manifolds with the unperturbed ones, see Proposition 4.8. Finally, in Section 5 we complete the proof of Theorem 1.2 by setting up an equation for the difference $(\Delta y,\Delta z)(x,\epsilon)$ for $x\in i\mathbb{R}$. We identify this equation as a slow-fast system and derive a Fenichel-type normal form. From this normal form, we complete the proof of Theorem 1.2, see Section 5.2.
2. The $\epsilon=0$-system: The unperturbed invariant manifolds
In this section, we consider the unperturbed system:
$x^{\prime}$ $=x^{2}+a(y^{2}+z^{2})+F(x,y,z,0),$ (2.1)
$y^{\prime}$ $=-bxy+z+G(x,y,z,0),$
$z^{\prime}$ $=-bxz-y+H(x,y,z,0).$
obtained from (1.1) by setting $\epsilon=0$ (or equivalently from (1.1) by setting $\mu=\nu=0$). Let $S^{\pm}(\delta,\chi)\subset\mathbb{C}$ denote the local open sectors, centered along the positive (++) (negative $($-), respectively) axis, of radius $\delta>0$ and opening $\pi+\chi\in(\pi,2\pi)$:
$S^{\pm}(\delta,\chi)=\left\{x\in\mathbb{C}\,:\,0<|x|<\delta\,\mbox{ and }\,|\operatorname{Arg}(\pm x)|<\frac{\pi+\chi}{2}\right\},$
respectively, see Fig. 3(a). Notice that $0<\chi<\pi$.
Lemma 2.1.
Consider (2.1). Then for $\delta>0$ sufficiently small, there exist two unique local invariant manifolds of (2.1) as graphs over $S^{\pm}=S^{\pm}(\delta,\chi)$:
$(y,z)=\psi_{0}^{\pm}(x),\quad x\in S^{\pm}.$ (2.2)
Here $\psi_{0}^{\pm}:S^{\pm}\rightarrow\mathbb{C}^{2}$ are each real-analytic and $\psi_{0}^{\pm}(x)=\mathcal{O}(x^{3})$ uniformly on the closure of $S^{\pm}$. In particular, they are the $1$-sums of a Gevrey-$1$ series:
$\psi_{0}^{\pm}(x)\sim_{1}\sum_{\alpha=3}^{\infty}\psi_{0,\alpha}x^{\alpha},\quad|\psi_{0,\alpha}|\leq c_{1}c_{2}^{\alpha}\alpha!\quad\forall\,\alpha\geq 3,$ (2.3)
for some $c_{1},c_{2}>0$, in the directions defined by $S^{\pm}$, respectively. In particular, the series in (2.3) is independent of $\pm$.
Proof.
We define $(r_{1},y_{1},z_{1})$ by
$\begin{cases}x=r_{1},\\ y=r_{1}^{2}y_{1},\\ z=r_{1}^{2}z_{1}.\end{cases}$ (2.4)
This gives
$r_{1}^{\prime}$ $=r_{1}^{2}(1+\mathcal{O}(r_{1})),$ (2.5)
$\zeta_{1}^{\prime}$ $=\Omega\zeta_{1}+\mathcal{O}(r_{1}),$
where
$\Omega:=\begin{pmatrix}0&1\\ -1&0\end{pmatrix}$, (2.6)
upon setting
$\zeta_{1}=(y_{1},z_{1})^{\mathrm{T}},$
or as a first order real-analytic system
$r_{1}^{2}\frac{d\zeta_{1}}{dr_{1}}$ $=\Omega\zeta_{1}+\mathcal{O}(r_{1}).$
This system is a generalized saddle-node with Poincaré rank $1$ studied by many authors, see e.g. [8]. In particular, the linearization of the right hand side for $r_{1}=0$ at $\zeta_{1}=(0,0)$ is given by $\Omega$, having nonzero eigenvalues $\pm i$. It then follows from [8, Theorem 3] (as a corollary of the invariance of $y=0$ for the normal form [8, Eq. (8)], see also [29, Lemma 4.8] or [16, Corollary 4.2] where the statement is made more explicit) that there exists two solutions
$\zeta_{1}=\psi^{\pm}_{1,0}(r_{1}),\quad r_{1}\in S^{\pm},$ (2.7)
with $\psi_{1,0}^{\pm}(0)=0$ and $S^{\pm}$ as in the statement. In particular, $\psi^{\pm}_{1,0}:S^{\pm}\rightarrow\mathbb{C}^{2}$ are each real-analytic, the $1$-sum of a Gevrey-$1$ series, with $\psi^{\pm}_{1,0}(r_{1})=\mathcal{O}(r_{1})$ uniformly on the closure of $S^{\pm}$, respectively. We complete the theorem by setting $\psi_{0}^{\pm}(r_{1}):=r_{1}^{2}\psi^{\pm}_{1,0}(r_{1})$, cf. (2.4). ∎
It is clear that $(y,z)=\psi_{0}^{\pm}(x)$, $x\in S^{\pm}\cap\mathbb{R}$, define unstable/stable sets, respectively, of the zero-Hopf point $(x,y,z)=(0,0,0)$ for $\epsilon=0$, see Fig. 3(b).
We refer to the invariant manifolds $(y,z)=\psi^{\pm}_{0}(x)$, $x\in S^{\pm}$, of (1.5), as the unperturbed manifolds.
Figure 3. In (a): The sectors $S^{\pm}$ in the complex plane. In (b): The invariant graphs (2.2) of (1.5) for $\epsilon=0$ in the real $(x,y,z)$-space.
3. The perturbed system: The unstable and stable manifolds
In this section, we consider (1.9) for all $0<\epsilon\ll 1$, repeated here in the fast form
$x_{2}^{\prime}$ $=\epsilon\left(x_{2}^{2}-1+a\epsilon^{2}(y_{2}^{2}+z_{2}^{2})+\epsilon F_{2}(x_{2},\epsilon y_{2},\epsilon z_{2},\epsilon)\right),$ (3.1)
$y_{2}^{\prime}$ $=\epsilon b(-x_{2}+\sigma)y_{2}+z_{2}+\epsilon G_{2}(x_{2},\epsilon y_{2},\epsilon z_{2},\epsilon),$
$z_{2}^{\prime}$ $=\epsilon b(-x_{2}+\sigma)z_{2}-y_{2}+\epsilon H_{2}(x_{2},\epsilon y_{2},\epsilon z_{2},\epsilon),$
with $(\cdot)^{\prime}=\frac{d}{d\tau}$ and where $F_{2},G_{2},H_{2}$ are defined by
$W_{2}(x_{2},\epsilon y_{2},\epsilon z_{2},\epsilon):=\epsilon^{-3}W(\epsilon x_{2},\epsilon(\epsilon y_{2}),\epsilon(\epsilon z_{2}),\epsilon),\quad W=F,G,H.$
Notice that $F_{2},G_{2}$ and $H_{2}$ have analytic extensions to $\epsilon=0$ by (1.6). In particular, as $F$, $G$ and $H$ are assumed to be real-analytic on a neighborhood of $(x,y,z,\mu)=(0,0,0,0)$, it follows that $F_{2}$, $G_{2}$ and $H_{2}$ are analytic on a polydisc:
$(x_{2},\epsilon y_{2},\epsilon z_{2},\epsilon)\in(\epsilon^{-1}B_{\xi})\times(\epsilon^{-1}B_{\xi})\times(\epsilon^{-1}B_{\xi})\times B_{\xi},$ (3.2)
for $\xi>0$ small enough. Here $B_{\xi}\subset\mathbb{C}$ denotes the open disc of radius $\xi>0$ centered at the origin. We recall that (3.1) is obtained from (1.5) by the change of coordinates
$(x_{2},y_{2},z_{2})\mapsto\begin{cases}x=\epsilon x_{2},\\ y=\epsilon^{2}y_{2},\\ z=\epsilon^{2}z_{2}.\end{cases}$ (3.3)
The system (1.9) is a slow-fast system for $0<\epsilon\ll 1$, with $x_{2}$ being slow and $(y_{2},z_{2})$ being fast. For $\epsilon=0$, we obtain the layer problem
$x_{2}^{\prime}$ $=0,$ (3.4)
$y_{2}^{\prime}$ $=z_{2},$
$z_{2}^{\prime}$ $=-y_{2},$
with $(x_{2},0,0)$ defining a critical manifold, which is normally elliptic. In terms of the slow time $t=\epsilon\tau$, we have
$\dot{x}_{2}$ $=x_{2}^{2}-1+a\epsilon^{2}(y_{2}^{2}+z_{2}^{2})+\epsilon F_{2}(x_{2},\epsilon y_{2},\epsilon z_{2},\epsilon),$ (3.5)
$\epsilon\dot{y}_{2}$ $=\epsilon b(-x_{2}+\sigma)y_{2}+z_{2}+\epsilon G_{2}(x_{2},\epsilon y_{2},\epsilon z_{2},\epsilon),$
$\epsilon\dot{z}_{2}$ $=\epsilon b(-x_{2}+\sigma)z_{2}-y_{2}+\epsilon H_{2}(x_{2},\epsilon y_{2},\epsilon z_{2},\epsilon),$
as in (1.9), which for $\epsilon=0$ gives the reduced problem:
$\dot{x}_{2}=Q_{2}(x_{2}):=x_{2}^{2}-1,$ (3.6)
for $y_{2}=z_{2}=0$. We have hyperbolic singularities at $x_{2}=\pm 1$, being unstable and stable, respectively. We illustrate the dynamics of (3.6), $x_{2}\in\mathbb{C}$, on the Poincaré sphere [17] in Fig. 4. The compactification is based upon
$\begin{cases}\operatorname{Re}(x_{2})={\breve{u}}{\breve{w}}^{-1},\\ \operatorname{Im}(x_{2})={\breve{v}}{\breve{w}}^{-1},\end{cases}$
with $(\breve{u},\breve{v},\breve{w})\in\mathbb{H}^{2}$. Here $\mathbb{H}^{2}\subset\mathbb{R}^{3}$ denotes the hemisphere:
$\mathbb{H}^{2}:=\{(\breve{u},\breve{v},\breve{w})\in\mathbb{R}^{3}\,:\,\breve{u}^{2}+\breve{v}^{2}+\breve{w}^{2}=1,\,\breve{w}\geq 0\}.$
The equator circle $\mathbb{S}^{1}=\mathbb{H}^{2}\cap\{\breve{w}=0\}$ represents $x_{2}=\infty$. There is a one-to-one correspondence between $x_{2}\in\mathbb{C}$ and $(\breve{u},\breve{v},\breve{w})\in\operatorname{int}\mathbb{H}^{2}:=\mathbb{H}^{2}\cap\{\breve{w}>0\}$. As usual in Poincaré compactification, we use a desingularization corresponding to division of the right hand side by $1+|x_{2}|^{2}$ to ensure that the flow is well-defined on the equator circle.
Figure 4. Phase portrait of $x_{2}^{\prime}=x_{2}^{2}-1$ on the Poincaré sphere (here represented as a disc). The boundary circle $\mathbb{S}^{1}$ represents $x_{2}=\infty$. The compact domains $\mathcal{X}^{\mp}$, used in Proposition 3.3, are subsets of the basin of attraction of $x_{2}=\mp 1$ for the forward and backward flow, respectively.
Lemma 3.1.
Consider (3.1) with $0<\epsilon\ll 1$ and suppose (1.12). Then there exist two hyperbolic saddle-foci $E^{\pm}(\epsilon)$ given by
$(x_{2},y_{2},z_{2})=(X^{\pm}({\epsilon}),\epsilon Y^{\pm}(\epsilon),\epsilon Z^{\pm}({\epsilon})),\quad X^{\pm}(0)=\pm 1,$
for all $0<\epsilon<\epsilon_{0}$, $0<\epsilon_{0}\ll 1$, respectively. Here $X^{\pm},Y^{\pm},Z^{\pm}$ are all real-analytic with respect to ${\epsilon}\in[0,\epsilon_{0})$. Moreover, $E^{+}(\epsilon)$ has a one-dimensional unstable manifold $\mathbf{W}^{u}(E^{+}(\epsilon))$ and a two-dimensional stable manifold whereas $E^{-}(\epsilon)$ has a one-dimensional stable manifold $\mathbf{W}^{s}(E^{-}(\epsilon))$ and a two-dimensional unstable manifold.
Proof.
Singularities of (3.1) are given by
$0$ $=x_{2}^{2}-1+a\epsilon^{2}(y_{2}^{2}+z_{2}^{2})+\epsilon F_{2}(x_{2},\epsilon y_{2},\epsilon z_{2},\epsilon),$
$0$ $=\epsilon b(-x_{2}+\sigma)y_{2}+z_{2}+\epsilon G_{2}(x_{2},\epsilon y_{2},\epsilon z_{2},\epsilon),$
$0$ $=\epsilon b(-x_{2}+\sigma)z_{2}-y_{2}+\epsilon H_{2}(x_{2},\epsilon y_{2},\epsilon z_{2},\epsilon).$
The existence and expansion of $E^{\pm}$ then follow from the implicit function theorem, using the regular solutions $(x_{2},y_{2},z_{2})=(\pm 1,0,0)$ for $\epsilon=0$. Subsequently, the statements regarding stability follow from simple calculations using (1.12). ∎
We illustrate the results of Lemma 3.1 in Fig. 5.
Figure 5. Illustration of the results of Lemma 3.1. In (a): $\epsilon=0$. In (b): $0<\epsilon\ll 1$.
In the following, we describe the local invariant manifolds $\mathbf{W}_{loc}^{s}(E^{-}(\epsilon))$ and $\mathbf{W}_{loc}^{u}(E^{+}(\epsilon))$ as graphs $(y_{2},z_{2})=m_{2}^{\pm}(x_{2})$ over $x_{2}$. These manifolds are solutions of the invariance equation:
$\left(\Omega+\epsilon b(-x_{2}+\sigma)\operatorname{Id}\right)\begin{pmatrix}y_{2}\\ z_{2}\end{pmatrix}+\epsilon\begin{pmatrix}G_{2}\\ H_{2}\end{pmatrix}=\epsilon\left(x_{2}^{2}-1+a\epsilon^{2}(y_{2}^{2}+z_{2}^{2})+\epsilon F_{2}\right)\frac{d}{dx_{2}}\begin{pmatrix}y_{2}\\ z_{2}\end{pmatrix}$, (3.7)
where $W_{2}=W_{2}(x_{2},y_{2},z_{2},\epsilon)$, $W=F,G,H$, and
$\operatorname{Id}=\operatorname{diag}(1,1)\in\mathbb{C}^{2\times 2}.$
Now, we recall the following:
Lemma 3.2.
There exists a unique formal series solution of (3.7) of the form
$(y_{2},z_{2})=\sum_{\alpha=1}^{\infty}m_{2,\alpha}(x_{2})\epsilon^{\alpha},$ (3.8)
with $m_{2,\alpha}:\mathbb{C}\rightarrow\mathbb{C}^{2}$ real-analytic for all $\alpha\in\mathbb{N}$.
Proof.
By (3.2), we have that for any $\varrho>0$, there is an $\epsilon_{0}>0$ such that $F_{2}$, $G_{2}$ and $H_{2}$ are real-analytic on $(x_{2},y_{2},z_{2},\epsilon)\in B_{\varrho}^{3}\times(-\epsilon_{0},\epsilon_{0})$. Here $B_{\varrho}^{3}\subset\mathbb{C}^{3}$ is the ball of radius $\varrho>0$ centered at the origin. But then [15, Proposition 2.1] gives the existence and uniqueness of the formal series (3.8) with $m_{2,\alpha}:B_{\varrho}\rightarrow\mathbb{C}^{2}$ real-analytic for all $\alpha\in\mathbb{N}$ and any $\varrho>0$. The result therefore follows. ∎
We now fix any compact neighborhood $\mathcal{X}_{2}^{-}\subset\mathbb{C}$ of $x_{2}=-1$ within its basin of attraction for the forward flow of (3.6). We similarly fix any compact neighborhood $\mathcal{X}_{2}^{+}\subset\mathbb{C}$ of $x_{2}=1$ within its basin of attraction for the backward flow of (3.6). See Fig. 4. Then we have the following:
Proposition 3.3.
There is an $\epsilon_{0}>0$ small enough, such that the local unstable and stable manifolds, $\mathbf{W}_{\mathrm{loc}}^{u}(E^{+}(\epsilon))$ and $\mathbf{W}_{\mathrm{loc}}^{s}(E^{-}(\epsilon))$, are graphs over $x_{2}\in\mathcal{X}_{2}^{\pm}$, respectively:
$(y_{2},z_{2})$ $=m^{\pm}_{2}(x_{2},\epsilon),\quad x_{2}\in\mathcal{X}_{2}^{\pm},$
with $\epsilon(Y_{2}^{\pm},Z_{2}^{\pm})=m_{2}^{\pm}(X_{2}^{\pm},\epsilon)$, for any $\epsilon\in(0,\epsilon_{0})$. Each $m_{2}^{\pm}(\cdot,\epsilon):\mathcal{X}_{2}^{\pm}\rightarrow\mathbb{C}^{2}$ is real-analytic and $C^{\infty}$ with respect to $\epsilon\in[0,\epsilon_{0})$. In particular, the functions $m_{2}^{\pm}$ are both asymptotic to the formal series (3.8) as $\epsilon\rightarrow 0$:
$m^{\pm}_{2}(x_{2},\epsilon)\sim\sum_{\alpha=1}^{\infty}m_{2,\alpha}(x_{2})\epsilon^{\alpha},$ (3.9)
uniformly with respect to $x_{2}\in\mathcal{X}_{2}^{\pm}$.
Proof.
The graph form of the invariant manifolds follows from standard hyperbolic theory, but due to the singular nature of $\epsilon\to 0$, this only gives local and nonuniform statements. Nevertheless, to prove the statement, we will follow the standard proof of the stable manifold theorem through the variation of constant formulation, after some initial preparation of the equations. We focus on $m_{2}^{-}$ and the stable manifold of $E^{-}$. The case of $m_{2}^{+}$ is similar and details of this case are therefore left out.
To prepare the equations, we first define
$\begin{cases}x_{2}=:X_{2}^{-}(\epsilon)+\widetilde{x}_{2},\\ y_{2}=:\epsilon Y_{2}^{-}(\epsilon)+\widetilde{y}_{2},\\ z_{2}=:\epsilon Z_{2}^{-}(\epsilon)+\widetilde{z}_{2},\end{cases}$
recall Lemma 3.1, and
$\zeta_{2}:=\widetilde{x}^{-1}\begin{pmatrix}\widetilde{y}_{2}\\ \widetilde{z}_{2}\end{pmatrix}$.
A simple calculation shows that $(\widetilde{x}_{2},\zeta_{2})\mapsto(x_{2},y_{2},z_{2})$ brings (3.1) into the following form by pull-back
$\dot{\widetilde{x}}_{2}$ $=\widetilde{Q}_{2}(\widetilde{x}_{2},\epsilon)\widetilde{x}_{2},$ (3.10)
$\epsilon\dot{\zeta}_{2}$ $=\left(\Omega+\epsilon W_{21}(\widetilde{x}_{2},\epsilon)\right)\zeta_{2}+\epsilon W_{20}(\widetilde{x}_{2},\epsilon)+\epsilon^{2}W_{22}(\widetilde{x}_{2},\zeta_{2},\epsilon)(\zeta_{2},\zeta_{2}),$
recall (2.6), after dividing the right hand side by a real-analytic quantity $1+\mathcal{O}(\epsilon^{2})$, uniformly with respect to $1+\widetilde{x}_{2}\in\mathcal{X}_{2}^{-}$, $|\zeta_{2}|\leq c$, with $c>0$ fixed, and all $0\leq\epsilon\ll 1$; notice that such division is justified by the fact that we are interested in invariant manifold solutions $\zeta_{2}=\zeta_{2}(\widetilde{x}_{2},\epsilon)$, satisfying the corresponding invariance equation:
$\left(\Omega+\epsilon W_{21}(\widetilde{x}_{2},\epsilon)\right)\zeta_{2}+\epsilon W_{20}(\widetilde{x}_{2},\epsilon)+\epsilon^{2}W_{22}(\widetilde{x}_{2},\zeta_{2},\epsilon)(\zeta_{2},\zeta_{2})=\epsilon\widetilde{Q}_{2}(\widetilde{x}_{2},\epsilon)\widetilde{x}_{2}\frac{\partial\zeta_{2}}{\partial\widetilde{x}_{2}}.$ (3.11)
In the derivation of (3.10), we have Taylor-expanded the right hand side of the $\zeta_{2}$-equation around $\zeta_{2}=0$ and used that $Q_{2}(x_{2}):=x_{2}^{2}-1\neq 0$ for all $x_{2}\in\mathcal{X}_{2}^{-}\setminus\{-1\}$ and that $Q_{2}^{\prime}(-1)=-2\neq 0$.
All quantities in (3.10) are real-analytic. In particular, we have
$\widetilde{Q}_{2}(\widetilde{x}_{2},0)=\widetilde{x}_{2}^{-1}Q_{2}(-1+\widetilde{x}_{2})=-2+\widetilde{x}_{2},$
and
$W_{21}(\widetilde{x}_{2},0)=\left(b(1-\widetilde{x}_{2}+\sigma)+2-\widetilde{x}_{2}\right)\operatorname{Id},$
specifically
$W_{21}(0,0)=(b(1+\sigma)+2)\operatorname{Id},$ (3.12)
with $b(1+\sigma)+2>0$. Finally, $W_{22}(\widetilde{x}_{2},\zeta_{2},\epsilon)(\cdot,\cdot)$ is bilinear.
As in Lemma 3.2, we have a unique formal series solution of (3.11) of the form
$\zeta_{2}=\sum_{\alpha=1}^{\infty}\zeta_{2,\alpha}(\widetilde{x}_{2})\epsilon^{\alpha},$ (3.13)
with $\zeta_{2,\alpha}:\mathbb{C}\rightarrow\mathbb{C}$ real-analytic for all $\alpha\in\mathbb{N}$. By uniqueness of the formal series, (3.13) corresponds to (3.8) upon the change of coordinates. We therefore fix $N\in\mathbb{N}$ and introduce ${\widetilde{\zeta}}_{2}$ by
$\zeta_{2}=\sum_{\alpha=1}^{2N}\zeta_{2,\alpha}(\widetilde{x}_{2})\epsilon^{\alpha}+\epsilon^{N}\widetilde{\zeta}_{2}.$ (3.14)
This brings the equations into the final form
$\dot{\widetilde{x}}_{2}$ $=\widetilde{Q}_{2}(\widetilde{x}_{2},\epsilon)\widetilde{x}_{2},$ (3.15)
$\epsilon\dot{\widetilde{\zeta}}_{2}$ $=\left(\Omega+\epsilon\widetilde{W}_{21}^{N}(\widetilde{x}_{2},\epsilon)\right)\widetilde{\zeta}_{2}+\epsilon^{N}\widetilde{W}_{20}^{N}(\widetilde{x}_{2},\widetilde{\zeta}_{2},\epsilon),$
with $\widetilde{W}_{21}^{N}(0,0)=(b(1+\sigma)+2)\operatorname{Id}$. Here $\widetilde{W}_{21}^{N}$ and $\widetilde{W}_{20}^{N}$ are real-analytic. This follows from a simple calculation.
We will prove the statement by working on (3.15). In particular, we define $\widetilde{\mathcal{X}}_{2}^{-}\subset\mathbb{C}$ by $\widetilde{x}_{2}\in\widetilde{\mathcal{X}}_{2}^{-}$ if and only if $1+\widetilde{x}_{2}\in\mathcal{X}_{2}^{-}$ and look for invariant manifold solutions $\widetilde{\zeta}_{2}=\widetilde{\zeta}_{2}(\widetilde{x}_{2},\epsilon)$ as graphs over $\widetilde{x}_{2}\in\mathcal{X}_{2}^{-}$.
We therefore take $\widetilde{x}_{20}\in\widetilde{\mathcal{X}}_{2}^{-}$ and consider the solution $\widetilde{x}_{2}(t,\widetilde{x}_{20},\epsilon)$ of the initial value problem
$\dot{\widetilde{x}}_{2}(t)=\widetilde{Q}_{2}(\widetilde{x}_{2}(t),\epsilon)\widetilde{x}_{2}(t),\quad\widetilde{x}_{2}(0)=\widetilde{x}_{20}.$
By assumption, $\mathcal{X}_{2}^{-}$ is compact and a subset of the basin of attraction of the hyperbolic point $x_{2}=-1$ for the forward flow of (3.6) for $\epsilon=0$. We therefore conclude that $\widetilde{x}_{2}(t,\widetilde{x}_{20},\epsilon)\rightarrow 0$ for $t\rightarrow\infty$, uniformly for $\widetilde{x}_{20}\in\mathcal{X}^{-}$, $0\leq\epsilon\ll 1$.
We then solve for $\widetilde{\zeta}_{2}(t,\widetilde{x}_{20},\epsilon)$, $t\geq 0$, such that $(\widetilde{x}_{20},\widetilde{\zeta}_{2}(0,\widetilde{x}_{20},\epsilon))$ is a point on the stable manifold through the usual fixed-point formulation of (3.15):
$\widetilde{\zeta}_{2}(t,\widetilde{x}_{20},\epsilon)=-\int_{t}^{\infty}\widetilde{\mathcal{M}}_{2}^{N}(t,\widetilde{x}_{20},\epsilon))\widetilde{\mathcal{M}}_{2}^{N}(s,\widetilde{x}_{20},\epsilon)^{-1}\epsilon^{N}\widetilde{W}_{20}^{N}(\widetilde{x}_{2}(s,\widetilde{x}_{20},\epsilon),\widetilde{\zeta}_{2}(s,\widetilde{x}_{20},\epsilon),\epsilon)ds,$ (3.16)
for $t\geq 0$, see e.g. [38, Theorem 5.3]. Here $\widetilde{\mathcal{M}}_{2}^{N}(t,\widetilde{x}_{20},\epsilon)$ is the fundamental matrix associated with the linear problem
$\epsilon\dot{\widetilde{\zeta}}_{2}(t)$ $=\left(\Omega+\epsilon\widetilde{W}_{21}^{N}(\widetilde{x}_{2}(t,\widetilde{x}_{20},\epsilon),\epsilon)\right)\widetilde{\zeta}_{2}(t),$
with $\widetilde{\mathcal{M}}_{2}^{N}(0,\widetilde{x}_{20},\epsilon)=\operatorname{Id}$, which by (3.12) has exponential growth for $t\to\infty$. Writing
$\widetilde{\mathcal{M}}_{2}^{N}(t,\widetilde{x}_{20},\epsilon)=:\exp({\epsilon^{-1}\Omega t})\widehat{\mathcal{M}}_{2}^{N}(t,\widetilde{x}_{20},\epsilon),$ (3.17)
we see that $\widehat{\mathcal{M}}_{N}(t,\widetilde{x}_{20},\epsilon)$ is the fundamental matrix associated with the regular problem:
$\dot{\widehat{\zeta}}_{2}(t)$ $=\widetilde{W}_{21}^{N}(\widetilde{x}_{2}(t,\widetilde{x}_{20},\epsilon),\epsilon)\widehat{\zeta}_{2}(t),$
with $\widehat{\mathcal{M}}_{2}^{N}(0,\widetilde{x}_{20},\epsilon)=\operatorname{Id}$. By assumption, we have the exponential convergence for $t\to\infty$:
$\widetilde{W}_{21}^{N}(\widetilde{x}_{2}(t,\widetilde{x}_{20},0),0)\to(b(1+\sigma)+2))\operatorname{Id},$
cf. (3.12), and from this we directly obtain
$\|\widetilde{\mathcal{M}}_{2}^{N}(t,\widetilde{x}_{20},\epsilon)\widetilde{\mathcal{M}}_{2}^{N}(s,\widetilde{x}_{20},\epsilon)^{-1}\|\leq c_{1}^{-1}{\mathrm{e}}^{c_{2}(t-s)},$ (3.18)
see e.g. [13, Proposition 1, p. 34], for $c_{1}>0,c_{2}>0$ both small enough, for any $s\geq t$, $\widetilde{x}_{2}\in\widetilde{\mathcal{X}}_{2}^{-}$, and all $0<\epsilon\ll 1$. Here $\|\cdot\|$ denotes the operator norm. To obtain (3.18), we have also used that $\|\exp({\epsilon^{-1}\Omega t})\|\equiv 1$, recall (2.6).
The existence of a unique bounded solution $t\mapsto\widetilde{\zeta}_{2}(t,\widetilde{x}_{20},\epsilon),t\geq 0$, of the fixed-point equation (3.16), then proceeds completely analogously to [38, Theorem 5.3] for any $0<\epsilon\ll 1$. The desired graph is given by $\widetilde{\zeta}_{2}=\widetilde{\zeta}_{2}(0,\widetilde{x}_{2},\epsilon)=\mathcal{O}(\epsilon^{N})$, uniformly over $\widetilde{x}_{2}\in\widetilde{\mathcal{X}}_{2}^{-}$, for all $0\leq\epsilon\ll 1$. Next, for the smoothness with respect to $\epsilon$, we first note that the stable manifold is unique, being independent of $N$, and smooth (even analytic) for $\epsilon\in(0,\epsilon_{0})$ by standard theory. The boundedness of $\widetilde{\zeta}_{2}$ for any $N$ and $0<\epsilon\ll 1$ then implies (by (3.14)) that the invariant manifold has an asymptotic expansion with respect to $\epsilon\to 0$, uniformly with respect to $x_{2}\in\mathcal{X}_{2}^{-}$. Finally, for any $k\in\mathbb{N}$ there is an $N\in\mathbb{N}$ ($N\geq 2k+1$ will do) such that the right hand side of (3.16) is $C^{k}$-smooth with respect to $\epsilon\in[0,\epsilon_{0})$; notice that we lose two powers of $\epsilon$ upon differentiation due to (3.17). It follows that the invariant manifold is $C^{\infty}$-smooth with respect to $\epsilon\in[0,\epsilon_{0})$. ∎
4. Blowup
Since the unstable and stable manifolds have the same asymptotic expansions (cf. (3.9)), we cannot determine the separation of the invariant manifolds by working in compact subsets of $x_{2}\in\mathbb{C}$. Notice that $x_{2}=\mathcal{O}(1)$ correspond to $x=\mathcal{O}(\epsilon)$ as $\epsilon\to 0$ by (1.8). We therefore wish to extend the unstable and stable manifolds to $x=\mathcal{O}(1)$ and compare the manifolds with the unperturbed manifolds, recall Lemma 2.1. For this purpose, we view (3.3) as a scaling chart ($\breve{\epsilon}=1$):
$\breve{\epsilon}=1:\quad\begin{cases}x=r_{2}x_{2}\\ y=r_{2}^{2}y_{2},\\ z=r_{2}^{2}z_{2},\\ \epsilon=r_{2},\end{cases}$ (4.1)
associated with the blowup transformation
$r\geq 0,\,(\breve{x},\breve{y},\breve{z},\breve{\epsilon})\in\mathbb{S}^{3}\mapsto\begin{cases}x=r\breve{x}\\ y=r^{2}\breve{y}\\ z=r^{2}\breve{z},\\ \epsilon=r\breve{\epsilon}\end{cases}.$ (4.2)
We will therefore also augment $\dot{\epsilon}=0$ to (1.5). To perform the extension, we use the $\breve{x}=1$-chart with associated coordinates $(r_{1},y_{1},z_{1},\epsilon_{1})$:
$\breve{x}=1:\quad\begin{cases}x=r_{1},\\ y=r_{1}^{2}y_{1},\\ z=r_{1}^{2}z_{1},\\ \epsilon=r_{1}\epsilon_{1}.\end{cases}$ (4.3)
The change of coordinates between $(x_{2},y_{2},z_{2},r_{2})$, see (4.1), and $(r_{1},y_{1},z_{1},\epsilon_{1})$, see (4.3), is well-defined for $x_{2}\neq 0$, $\epsilon_{1}\neq 0$, and given by
$\begin{cases}r_{1}=r_{2}x_{2},\\ \epsilon_{1}=x_{2}^{-1},\\ y_{1}=x_{2}^{-2}y_{2},\\ z_{1}=x_{2}^{-2}z_{2}.\end{cases}$ (4.4)
The advantage of working with (4.3), and the coordinates $(r_{1},y_{1},z_{1},\epsilon_{1})$, is that both results of Lemma 2.1 and Proposition 3.3 are “visible” within compact subsets. In particular, Lemma 2.1 is visible within $\epsilon_{1}=0$, see also (2.7), whereas Proposition 3.3 is visible for $r_{1}=\mathcal{O}(\epsilon)$, $\epsilon_{1}=\mathcal{O}(1)$ (cf. (4.4)).
4.1. Analysis in the $\breve{x}=1$-chart
In the $\breve{x}=1$-chart, we apply the change of coordinates $(r_{1},y_{1},z_{1},\epsilon_{1})\mapsto(x,y,z,\epsilon)$, defined by (4.3), to the extended vector-field given by (1.5) and $\dot{\epsilon}=0$. This leads to the following system by pull-back
$\dot{r}_{1}$ $=r_{1}^{2}\left[1-\epsilon_{1}^{2}+ar_{1}^{2}(y_{1}^{2}+z_{1}^{2})+r_{1}F_{1}(r_{1},r_{1}y_{1},r_{1}z_{1},\epsilon_{1})\right],$ (4.5)
$\dot{y}_{1}$ $=r_{1}b(-1+\epsilon_{1}\sigma)y_{1}+z_{1}+r_{1}G_{1}(r_{1},r_{1}y_{1},r_{1}z_{1},\epsilon_{1})$
$-2r_{1}\left[1-\epsilon_{1}^{2}+ar_{1}^{2}(y_{1}^{2}+z_{1}^{2})+r_{1}F_{1}(r_{1},r_{1}y_{1},r_{1}z_{1},\epsilon_{1})\right]y_{1},$
$\dot{z}_{1}$ $=r_{1}b(-1+\epsilon_{1}\sigma)z_{1}-y_{1}+r_{1}H_{1}(r_{1},r_{1}y_{1},r_{1}z_{1},\epsilon_{1})$
$-2r_{1}\left[1-\epsilon_{1}^{2}+ar_{1}^{2}(y_{1}^{2}+z_{1}^{2})+r_{1}F_{1}(r_{1},r_{1}y_{1},r_{1}z_{1},\epsilon_{1})\right]z_{1},$
$\dot{\epsilon}_{1}$ $=-r_{1}\epsilon_{1}\left[1-\epsilon_{1}^{2}+ar_{1}^{2}(y_{1}^{2}+z_{1}^{2})+r_{1}F_{1}(r_{1},r_{1}y_{1},r_{1}z_{1},\epsilon_{1})\right],$
where $F_{1},G_{1},H_{1}$ are defined by
$W_{1}(r_{1},r_{1}y_{1},r_{1}z_{1},\epsilon_{1}):=r_{1}^{-3}W(r_{1},r_{1}r_{1}y_{1},r_{1}r_{1}z_{1},r_{1}\epsilon_{1}),\quad W=F,G,H,$
each having analytic extensions to $r_{1}=0$. Moreover, by (1.3) it follows that
$F_{1}(0,0,0,\epsilon_{1})\equiv F_{0}.$ (4.6)
Notice that within the invariant set $\epsilon_{1}=0$, (4.5) reduces to (2.5), possessing invariant manifolds of the graph forms
$(y_{1},z_{1})=\psi_{1,0}^{\pm}(r_{1}),$
defined over the sectors $r_{1}\in S^{\pm}$, recall Lemma 2.1. Within $r_{1}=0$, the line defined by $(y_{1},z_{1})=(0,0)$, $\epsilon_{1}\geq 0$, is an elliptic critical manifold (for $\epsilon_{1}>0$ it corresponds to the critical manifold of the scaling chart: $\breve{\epsilon}=1$, recall (3.4)).
Since we are interested in invariant manifolds, we can multiply the right hand side of (4.5) by the nonzero quantity
$\frac{1-\epsilon_{1}^{2}+r_{1}F_{0}}{\left[1-\epsilon_{1}^{2}+ar_{1}^{2}(y_{1}^{2}+z_{1}^{2})+r_{1}F_{1}(r_{1},r_{1}y_{1},r_{1}z_{1},\epsilon_{1})\right]}=:1+\widetilde{F}_{1}(r_{1},r_{1}y_{1},r_{1}z_{1},\epsilon_{1}),$
with
$\widetilde{F}_{1}(r_{1},r_{1}y_{1},r_{1}z_{1},\epsilon_{1})=\mathcal{O}(r_{1}^{2}).$ (4.7)
Here we have used (4.6). Then (4.5) becomes
$\dot{r}_{1}$ $=r_{1}^{2}\left(1-\epsilon_{1}^{2}+r_{1}F_{0}\right),$ (4.8)
$\dot{\zeta}_{1}$ $=\left(\Omega+r_{1}W_{11}(r_{1},\epsilon_{1})\right)\zeta_{1}+r_{1}W_{10}(r_{1},\epsilon_{1})+r_{1}^{2}W_{12}(r_{1},\zeta_{1},\epsilon_{1})(\zeta_{1},\zeta_{1}),$
$\dot{\epsilon}_{1}$ $=-r_{1}\epsilon_{1}\left(1-\epsilon_{1}^{2}+r_{1}F_{0}\right),$
upon setting $\zeta_{1}=(y_{1},z_{1})^{\mathrm{T}}$, for some real-analytic functions $W_{1i}$, $i\in\{0,1,2\}$. In particular, we have
$W_{11}(0,\epsilon_{1})=(-2-b-\epsilon_{1}\sigma)\operatorname{Id}.$
This follows from a simple calculation. In the derivation of (4.8), we have Taylor-expanded the right hand side of the $\zeta_{1}$-equation around $\zeta_{1}=0$. Notice, in particular, that
$W_{12}(r_{1},\zeta_{1},\epsilon_{1})(\cdot,\cdot),$
is bilinear.
With slight abuse of notation, we will use $\dot{(\cdot})=\frac{d}{dt}$ in the following.
We now look for (formal) invariant manifold solutions of (4.8) of the graph form $\zeta_{1}=\zeta_{1}(r_{1},\epsilon_{1})$ (which by virtue of the multiplication by the nonzero quantity (4.7) are also (formal) invariant manifolds of (4.5)). They satisfy the PDE
$r_{1}^{2}\left(1-\epsilon_{1}^{2}+r_{1}F_{0}\right)\frac{\partial\zeta_{1}}{\partial r_{1}}-r_{1}\epsilon_{1}\left(1-\epsilon_{1}^{2}+r_{1}F_{0}\right)\frac{\partial\zeta_{1}}{\partial\epsilon_{1}}$ (4.9)
$=\left(\Omega+r_{1}W_{11}(r_{1},\epsilon_{1})\right)\zeta_{1}+r_{1}W_{10}(r_{1},\epsilon_{1})+r_{1}^{2}W_{12}(r_{1},\zeta_{1},\epsilon_{1})(\zeta_{1},\zeta_{1}),$
in the formal sense. In particular, we will (as in [29, Section 4.3]) use two different notions of formal series: Formal series in $r_{1}$ with $\epsilon_{1}$-depending coefficients:
$\zeta_{1}=\sum_{\alpha=1}^{\infty}m_{1,\alpha}(\epsilon_{1})r_{1}^{\alpha},$ (4.10)
and formal series in $\epsilon_{1}$ with $r_{1}$-depending coefficients:
$\zeta_{1}=\sum_{\alpha=0}^{\infty}\psi_{1,\alpha}(r_{1})\epsilon_{1}^{\alpha}.$ (4.11)
Since (4.9) is analytic, we can again expand and understand (4.9) as an equation for formal series in both cases. In the case (4.10), the coefficients $m_{1,\alpha}$ will be analytic in a neighborhood $B_{\xi}\subset\mathbb{C}$, $\xi>0$, of $\epsilon_{1}=0$. Here $B_{\xi}$ denotes the open disc in $\mathbb{C}$ of radius $\xi>0$ centered at the origin. On the other hand, in the case (4.11) the coefficients $\psi_{1,\alpha}$ will come in pairs being analytic over the local sectors $r_{1}\in S^{\pm}$, respectively.
Lemma 4.1.
Fix $\xi>0$ small enough. Then there exists a unique formal series solution of (4.9) of the form (4.10) where each $m_{1,\alpha}:B_{\xi}\rightarrow\mathbb{C}^{2}$, $\alpha\in\mathbb{N}$, is a uniquely determined real-analytic function. Here $\xi>0$ is small enough, but independent of $\alpha$.
Proof.
We write (4.9) in the form
$\zeta_{1}=\Omega^{-1}r_{1}(1-\epsilon_{1}^{2})\left(r_{1}\frac{\partial\zeta_{1}}{\partial r_{1}}-\epsilon_{1}\frac{\partial\zeta_{1}}{\partial\epsilon_{1}}-R_{1}(r_{1},\zeta_{1},\epsilon_{1})\right),$ (4.12)
for some locally defined real-analytic function $R_{1}$. This follows from a simple calculation. The result can then be proven as the proof of [15, Proposition 2.1.] (by the virtue of the $r_{1}$-factor on the right hand side). ∎
Remark 4.2.
Upon using the change of coordinates (4.4), (4.10) becomes
$\zeta_{2}$ $=\sum_{\alpha=1}^{\infty}m_{1,\alpha}(x_{2}^{-1})x_{2}^{\alpha+2}r_{2}^{\alpha},$
and therefore by the uniqueness of the series, we have that
$m_{2,\alpha}(x_{2})=m_{1,\alpha}(x_{2}^{-1})x_{2}^{\alpha{+}2},\quad x_{2}\neq 0,$
for all $\alpha\in\mathbb{N}$, see (3.9).
We recall the definition of $S^{\pm}(\delta,\chi)$:
$S^{\pm}(\delta,\chi)=\left\{x\in\mathbb{C}\,:\,0<|x|<\delta\,\mbox{ and }\,|\operatorname{Arg}(\pm x)|<\frac{\pi+\chi}{2}\right\},$
with $0<\chi<\pi$ and $\delta>0$.
Lemma 4.3.
Fix $\delta>0$ small enough. Then there exists two (corresponding to $\pm$) formal series solution of (4.9) of the form (4.11) with $\psi_{1,\alpha}=\psi_{1,\alpha}^{\pm}$:
$\zeta_{1}=\sum_{\alpha=0}^{\infty}\psi^{\pm}_{1,\alpha}(r_{1})\epsilon_{1}^{\alpha},$ (4.13)
where each $\psi^{\pm}_{1,\alpha}:S^{\pm}\rightarrow\mathbb{C}$, $\alpha\in\mathbb{N}_{0}$, is real-analytic on the local sector
$S^{\pm}=S^{\pm}(\delta,\chi),$
respectively. Moreover, the two series are uniquely determined and for each fixed $\alpha\in\mathbb{N}_{0}$, $\psi^{\pm}_{1,\alpha}$ is the $1$-sum of a Gevrey-$1$ series
$\psi^{\pm}_{1,\alpha}(r_{1})\sim_{1}\sum_{\beta=1}^{\infty}\psi_{1,\alpha,\beta}r_{1}^{\beta},$
in the direction defined by $S^{\pm}$, respectively. In particular, the Gevrey-$1$ series on the right hand side is independent of $\pm$.
Proof.
We rearrange (4.12) to the following form
$r_{1}^{2}\frac{\partial\zeta_{1}}{\partial r_{1}}-r_{1}\epsilon_{1}\frac{\partial\zeta_{1}}{\partial\epsilon_{1}}=\Omega(1-\epsilon_{1}^{2})^{-1}\zeta_{1}+r_{1}R_{1}(r_{1},\zeta_{1},\epsilon_{1}).$ (4.14)
With $\zeta_{1}$ as the formal series (4.13) (for simplicity we drop the $\pm$-superscript, which refers to the direction $S^{\pm}$), we use the Faa di Bruno formula [12, Theorem 2.1] (with $F(\zeta_{1},\epsilon_{1})=R_{1}(r_{1},\zeta_{1},\epsilon_{1})$, $G(\epsilon_{1})=(\zeta_{1}(\epsilon_{1}),\epsilon_{1})$) to write
$R_{1}(r_{1},\zeta_{1},\epsilon_{1})$ $=R_{1}(r_{1},\psi_{1,0},0)+\sum_{\alpha=1}^{\infty}R_{1,\alpha}(r_{1},\psi_{1,0},\ldots,\psi_{1,\alpha})\epsilon_{1}^{\alpha},$ (4.15)
where
$R_{1,\alpha}(r_{1},\psi_{1,0},\ldots,\psi_{1,\alpha})=$ $\sum_{1\leq|\lambda|\leq\alpha}\frac{1}{\lambda!}\frac{\partial^{|\lambda|}}{\partial\zeta_{1}^{(\lambda_{1},\lambda_{2})}\partial\epsilon_{1}^{\lambda_{3}}}R_{1}(r_{1},\psi_{1,0},0)\times$ (4.16)
$\sum_{\gamma=1}^{\alpha}\sum_{(\sigma,\ell)\in p_{\gamma,\alpha,\lambda}}\prod_{\beta=1}^{\gamma}\frac{1}{\sigma_{\beta}!(\ell_{\beta}!)^{|\sigma_{\beta}|}}(\ell_{\beta}!\psi_{1,\ell_{\beta}})^{(\sigma_{1\beta},\sigma_{2\beta})}(\delta_{1\beta})^{\sigma_{3\beta}}.$
Here $\delta_{\alpha\beta}$ denotes Kronecker’s delta. The details of the finite index sets
$p_{\gamma,\alpha,\lambda}\subset\left\{(\sigma,\ell)\in(\mathbb{N}^{3})^{\gamma}\times\mathbb{N}^{\gamma}\,:\,\sum_{\beta=0}^{\gamma}\sigma_{\beta}=\lambda\,\,\text{and}\,\,\sum_{\beta=0}^{\gamma}|\sigma_{\beta}|\ell_{\beta}=\alpha\right\}.$ (4.17)
are not important, and we therefore refer to [12, Eq. (2.2)]. We will only use that
$R_{1,\alpha}(r_{1},\psi_{1,0},\ldots,\psi_{1,\alpha}):=Q_{1,\alpha}(r_{1},\psi_{1,0},\ldots,\psi_{1,\alpha-1})+\frac{\partial}{\partial\zeta_{1}}R_{1}(r_{1},\psi_{1,0},0)\psi_{1,\alpha},$
with the last term obtained from the sum in (4.16) with $\lambda=(1,0,0)$ and $\lambda=(0,1,0)$ (where $p_{\gamma,\alpha,\lambda}$ are singletons), so that $Q_{1,\alpha}$ (cf. (4.16) and (4.17)) is a multivariate polynomial of degree $\alpha$ with respect to the variables
$(\psi_{1,1},\ldots,\psi_{1,\alpha-1}),$
for fixed $r_{1},\psi_{1,0}$, for all $\alpha\in\mathbb{N}$. In this way, by inserting (4.13) into (4.14) and collecting terms, we obtain the following equations for $\psi_{1,\alpha}$, $\alpha\in\mathbb{N}_{0}$:
$r_{1}^{2}\psi_{1,0}^{\prime}=\Omega\psi_{1,0}+r_{1}R_{1}(r_{1},\psi_{1,0},0),$
and
$r_{1}^{2}\psi_{1,\alpha}^{\prime}$ $=\Omega\psi_{1,\alpha}+r_{1}\left(\frac{\partial R_{1}}{\partial\zeta_{1}}(r_{1},\psi_{1,0},0)+\alpha\operatorname{Id}\right)\psi_{1,\alpha}$ (4.18)
$+\sum_{\gamma=0}^{\alpha-1}\Omega A_{1,\alpha-\gamma}\psi_{1,\gamma}+r_{1}Q_{1,\alpha-1}(r_{1},\psi_{1,0},\ldots,\psi_{1,\alpha-1})\quad\forall\,\alpha\in\mathbb{N},$
where
$A_{1,\alpha}=\begin{cases}1&\mbox{$\alpha$ even},\\ 0&\mbox{$\alpha$ odd}.\end{cases}$
It follows that $\psi_{1,\alpha}\in S^{\pm}(\delta_{\alpha},\chi)$ for $\delta_{\alpha}>0$ sufficiently small for all $\alpha\in\mathbb{N}_{0}$ from [8] (as in the proof of Lemma 2.1). In particular, $\psi_{1,\alpha}^{\pm}$ are the $1$-sums of a Gevrey-$1$ series in the directions defined by $S^{\pm}$, respectively. To obtain that $\delta_{\alpha}>0$ is uniformly bounded from below, we use that (4.18) is a linear first order equation for $\psi_{1,\alpha}$, $\alpha\in\mathbb{N}$, which is regular for $r_{1}\neq 0$. Therefore, by Lemma 2.1 ($\alpha=0$ case), the existence and uniqueness of solutions of linear equations and induction on $\alpha$, we conclude that $\psi_{1,\alpha}$ can be extended to $S^{\pm}(\delta,\chi)$, $\delta=\delta_{0}$, for all $\alpha\in\mathbb{N}_{0}$. ∎
We will now use the formal invariant solutions to bring (4.8) into a suitable normal form, where the unstable and stable manifolds from the scaling chart can be extended to $r_{1}=\mathcal{O}(1)$ through application of the forward flow.
In the following, we let $\mathcal{G}(S)\{\zeta_{1},\epsilon_{1}\}$ denote the set of analytic functions $W_{1}:S\times B_{\xi}^{3}\rightarrow\mathbb{C}^{\gamma}$, where $S=S^{\pm}(\delta,\chi)$ is an open sector and where $B_{\xi}^{3}\subset\mathbb{C}^{3}$ is the open ball in $\mathbb{C}^{3}$ of radius $\xi>0$ centered at the origin, such that
$W_{1}(r_{1},\zeta_{1},\epsilon_{1})=\sum_{\alpha,\beta}W_{1,\alpha,\beta}(r_{1})\zeta_{1}^{\alpha}\epsilon_{1}^{\beta},$
with $\alpha,\beta$ ranging over $\mathbb{N}_{0}^{2},\mathbb{N}_{0}$, respectively. Here the series is assumed to be absolutely convergent for $(\zeta_{1},\epsilon_{1})\in B_{\xi}^{3}\subset\mathbb{C}^{3}$, uniformly with respect $r_{1}\in S$, and where each $W_{1,\alpha,\beta}$ is the $1$-sum on $S$ (being continuous on the closure of $S$ with $W_{1,\alpha,\beta}(0)=0$ for all $\alpha\in\mathbb{N}_{0}^{2},\beta\in\mathbb{N}_{0}$) of a Gevrey-$1$ series as $r_{1}\rightarrow 0$. As indicated, we suppress $\gamma$, which should be clear from the context. $\mathcal{G}(S)\{\epsilon_{1}\}$ (as a set of functions that are independent of $\zeta_{1}$) is defined similarly.
In the following, we fix $\delta>0$ and $\xi>0$ small enough and set $S^{\pm}:=S^{\pm}(\delta,\chi)$ with $0<\chi<\pi$. Recall that $\epsilon=r_{1}\epsilon_{1}$.
Proposition 4.4.
Fix any $N\in\mathbb{N}$. Then there exist two functions $W_{1}^{\pm,N}\in\mathcal{G}(S^{\pm})\{\epsilon_{1}\}$ such that the (blowup) transformations
$(r_{1},\widetilde{\zeta}_{1},\epsilon_{1})\mapsto(r_{1},\zeta_{1},\epsilon_{1}),\quad r_{1}\in S^{\pm},\quad(\widetilde{\zeta}_{1},\epsilon_{1})\in B_{\xi}^{3},$ (4.19)
defined by
$\zeta_{1}=W_{1}^{\pm,N}(r_{1},\epsilon_{1})+(r_{1}\epsilon_{1})^{N}\widetilde{\zeta}_{1},$ (4.20)
bring (4.8) into the following two sets of equations by pull-back
$\dot{r}_{1}$ $=r_{1}^{2}\left(1-\epsilon_{1}^{2}+r_{1}F_{0}\right),$ (4.21)
$\dot{\widetilde{\zeta}}_{1}$ $=\left(\Omega+r_{1}\widetilde{W}_{11}^{\pm,N}(r_{1},\epsilon_{1})\right)\widetilde{\zeta}_{1}+(r_{1}\epsilon_{1})^{N}\widetilde{W}_{10}^{\pm,N}(r_{1},\widetilde{\zeta}_{1},\epsilon_{1}),$
$\dot{\epsilon}_{1}$ $=-r_{1}\epsilon_{1}\left(1-\epsilon_{1}^{2}+r_{1}F_{0}\right),$
for $r_{1}\in S^{\pm}$, $(\widetilde{\zeta}_{1},\epsilon_{1})\in B_{\xi}^{3}$. Here $\widetilde{W}_{11}^{\pm,N}\in\mathcal{G}(S^{\pm})\{\epsilon_{1}\}$, $\widetilde{W}_{10}^{\pm,N}\in\mathcal{G}(S^{\pm})\{\widetilde{\zeta}_{1},\epsilon_{1}\}$ and
$W_{1}^{\pm,N}(r_{1},0)=\psi_{1,0}^{\pm}(r_{1}),\quad W_{1}^{\pm,N}(0,\epsilon_{1})=0,$ (4.22)
for all $r_{1}\in S^{\pm}$, $\epsilon_{1}\in B_{\xi}$, respectively. Moreover,
$\widetilde{W}_{11}^{\pm,N}(0,\epsilon_{1})=(-2-b+\epsilon_{1}\sigma)\operatorname{Id}.$ (4.23)
Remark 4.5.
Notice that (4.19) is of blowup-type since it reduces to $\zeta_{1}=W_{1}^{\pm,N}(r_{1},\epsilon_{1})$ for any $(r_{1},\epsilon_{1})\,:\,r_{1}\epsilon_{1}=0$.
Proof.
We follow [29, Lemma 4.5] and take
$W_{1}^{\pm,N}(r_{1},\epsilon_{1}):=\sum_{\alpha=0}^{2N-1}\psi_{1,\alpha}^{\pm}(r_{1})\epsilon_{1}^{\alpha}+\sum_{\alpha=1}^{2N-1}(m_{1,\alpha}-J^{2N-1}(m_{1,\alpha}))(\epsilon_{1})r_{1}^{\alpha},$
[…]
$J^{\gamma}(W_{1})(\epsilon_{1}):=\sum_{\alpha=0}^{\gamma}W_{1,\alpha}\epsilon_{1}^{\alpha}.$
Clearly, $W_{1}^{\pm,N}\in\mathcal{G}(S^{\pm})\{\epsilon_{1}\}$, satisfying (4.22). Then by construction $\zeta_{1}=W_{1}^{\pm,N}(r_{1},\epsilon_{1})$ defines an invariant manifold up to error-terms of order $\mathcal{O}((r_{1}\epsilon_{1})^{2N})=\mathcal{O}(\epsilon^{2N})$, in the sense that there exists two locally defined real-analytic functions $\widetilde{R}_{1}^{\pm,N}\in\mathcal{G}(S^{\pm})\{\epsilon_{1}\}$:
$r_{1}^{2}\left(1-\epsilon_{1}^{2}+r_{1}F_{0}\right)\frac{\partial W_{1}^{\pm,N}}{\partial r_{1}}-r_{1}\epsilon_{1}\left(1-\epsilon_{1}^{2}+r_{1}F_{0}\right)\frac{\partial W_{1}^{\pm,N}}{\partial\epsilon_{1}}$
$-\left(\left(\Omega+r_{1}W_{11}\right)W_{1}^{\pm,N}+r_{1}W_{10}+r_{1}^{2}W_{12}(W_{1}^{\pm,N})^{2}\right)$
$=:(r_{1}\epsilon_{1})^{2N}\widetilde{R}_{1}^{\pm,N}(r_{1},\epsilon_{1})\quad\forall\,r_{1}\in S^{\pm},\,\epsilon_{1}\in B_{\xi},$
writing (for simplicity)
$W_{10}(r_{1},\epsilon_{1}),\,\,W_{11}(r_{1},\epsilon_{1}),\,\,W_{12}(r_{1},W_{1}^{\pm,N},\epsilon_{1})(W_{1}^{\pm,N},W_{1}^{\pm,N}),$
compactly as
$W_{10},\,\,W_{11},\,\,W_{12}(W_{1}^{\pm,N})^{2},\,\,\mbox{respectively}.$
Therefore (4.20) gives
$(r_{1}\epsilon_{1})^{N}\dot{\widetilde{\zeta}}_{1}$ $=\left(\Omega+r_{1}W_{11}\right)\zeta_{1}+r_{1}W_{10}+r_{1}^{2}W_{12}(\zeta_{1})^{2}$
$-r_{1}^{2}\left(1-\epsilon_{1}^{2}+r_{1}F_{0}\right)\frac{\partial W_{1}^{\pm,N}}{\partial r_{1}}$
$+r_{1}\epsilon_{1}\left(1-\epsilon_{1}^{2}+r_{1}F_{0}\right)\frac{\partial W_{1}^{\pm,N}}{\partial\epsilon_{1}}$
$=\left(\Omega+r_{1}W_{11}\right)(r_{1}\epsilon_{1})^{N}\widetilde{\zeta}_{1}-(r_{1}\epsilon_{1})^{2N}\widetilde{R}_{1}^{\pm,N}$
$+r_{1}^{2}\left(W_{12}\left(W_{1}^{\pm,N}+(r_{1}\epsilon_{1})^{N}\widetilde{\zeta}_{1})\right)^{2}-W_{12}\left(W_{1}^{\pm,N}\right)^{2}\right)$
$=:\left(\Omega+r_{1}\widetilde{W}_{11}^{\pm,N}(r_{1},\epsilon_{1})\right)(r_{1}\epsilon_{1})^{N}\widetilde{\zeta}_{1}+(r_{1}\epsilon_{1})^{2N}\widetilde{W}_{10}^{\pm,N}(r_{1},\widetilde{\zeta}_{1},\epsilon_{1}),$
where we have used a Taylor-expansion with respect to $\widetilde{\zeta}_{1}$ in the final equality. The expansion of $\widetilde{W}_{11}^{\pm,N}$ follows from a simple calculation using that $W_{1}^{\pm,N}(r_{1},\epsilon_{1})=\mathcal{O}(r_{1})$. ∎
Below we will use the flow of (4.21), $r_{1}\in S^{\pm}$, to extend the invariant manifolds from $x=r_{1}=\mathcal{O}(\epsilon)$ (recall Proposition 3.3) to $x=r_{1}=\mathcal{O}(1)$. We focus on the unstable manifold and the case corresponding to “++”. The case “-” can be handled in the same way and the details of this case is therefore left out.
4.2. Elliptic paths
In this section, we first focus on the $(r_{1},\epsilon_{1})$-subsystem of (4.21):
$\dot{r}_{1}$ $=r_{1}^{2}(1-\epsilon_{1}^{2}+r_{1}F_{0}),$ (4.24)
$\dot{\epsilon}_{1}$ $=-r_{1}\epsilon_{1}(1-\epsilon_{1}^{2}+r_{1}F_{0}),$
with $\dot{(\cdot})=\frac{d}{dt}$, which by construction decouples.
We call trajectories $(r_{1}(t),\epsilon_{1}(t)))$, $t\in I\subset\mathbb{R}$, elliptic paths (see e.g. [26]) since the $\widetilde{\zeta}_{1}$-dynamics, see (4.21), is dominated by oscillations along such paths, see also Lemma 4.7 below. In contrast to hyperbolic paths, see Section 5 below, where the $\widetilde{\zeta}_{1}$-dynamics is dominated by contraction and expansion, the elliptic paths offer a suitable control of the $\widetilde{\zeta}_{1}$-dynamics to extend the unstable manifold, see also [40, 41].
Notice that upon restricting attention to $\epsilon=r_{1}\epsilon_{1}=\textrm{const}.$, (4.24) can be reduced to
$\dot{r}_{1}$ $=r_{1}^{2}-\epsilon^{2}+r_{1}^{3}F_{0}.$ (4.25)
We focus on the closed region $\mathcal{A}^{+}(\epsilon)$ defined by $r_{1}\in S^{+}(\delta,\chi)$ and $|r_{1}|\in[\xi^{-1}\epsilon,\delta]$, $\xi>0$ and $\delta>0$ small enough, and all $0<\epsilon\ll 1$. In polar coordinates $r_{1}=\rho_{1}e^{i\phi}$, the region becomes $\rho_{1}\in[\xi^{-1}\epsilon,\delta]$, $\phi\in(-\frac{\pi+\chi}{2},\frac{\pi+\chi}{2})$. Moreover, in these coordinates (4.25) becomes
$\dot{\rho}_{1}$ $=\rho_{1}^{2}\left(\cos\phi-\rho_{1}^{-2}\epsilon^{2}\cos\phi+\rho_{1}\cos(2\phi)F_{0}\right),$ (4.26)
$\dot{\phi}$ $=\sin\phi\left(1+\rho_{1}^{-2}\epsilon^{2}+2\rho_{1}F_{0}\cos\phi\right)\rho_{1}.$
Clearly, $\dot{\phi}\gtrless 0$ for $\phi\gtrless 0$ whereas $\dot{\rho}_{1}=0$ for $\rho_{1}\in[\xi^{-1}\epsilon,\delta]$ gives two solutions
$\phi=\phi_{\pm}(\rho_{1},\epsilon^{2}\rho_{1}^{-2}):=\pm\frac{\pi}{2}+\mathcal{O}(\rho_{1},\epsilon^{2}\rho_{1}^{-2}),\quad\rho_{1}\in[\xi^{-1}\epsilon,\delta],$ (4.27)
within $\mathcal{A}^{+}(\epsilon)$. This latter property can be obtained from the implicit function theorem. Within the subset of $\mathcal{A}^{+}(\epsilon)$ defined by
$\phi\in(\phi_{-}(\rho_{1},\epsilon^{2}\rho_{1}^{-2}),\phi_{+}(\rho_{1},\epsilon^{2}\rho_{1}^{-2})),\quad\rho_{1}\in[\xi^{-1}\epsilon,\delta],$
we have $\rho_{1}^{\prime}>0$; recall that $\xi>0$ and $\delta>0$ are fixed small enough. We sketch the phase-portrait in Fig. 6.
Lemma 4.6.
Let $r_{1}(t)\in\mathcal{A}^{+}(\epsilon)$ denote a forward solution of (4.25) in $\mathcal{A}^{+}(\epsilon)$ with initial condition $r_{1}(0)$ in the interior of $\mathcal{A}^{+}(\epsilon)$. Then for any $0<\epsilon\ll 1$ there is a $T=T(r_{1}(0),\epsilon)>0$ such that $r_{1}(t)\in\mathcal{A}^{+}(\epsilon)$ for all $t\in[0,T)$ and $r_{1}(T)\in\partial\mathcal{A}^{+}(\epsilon)$. Moreover, there exists a constant $C>0$ such that
$T(r_{1}(0),\epsilon)\leq C\epsilon^{-1}\quad\forall\,r_{1}(0)\in\mathcal{A}^{+}(\epsilon),\,0<\epsilon\ll 1.$ (4.28)
Proof.
The existence of $T$ follows from the analysis above, see also Fig. 6 for an illustration. To obtain the bound, we consider (4.26). Suppose first that $\phi(t)\in[-c,c]$ with $c>0$ fixed small enough. Then
$\dot{\rho}_{1}>\frac{1}{2}\rho_{1}^{2}(1-\rho_{1}^{-2}\epsilon^{2}),$
for $\xi,\delta>0$ small enough, from which an upper bound of $T$ of the form $C\epsilon^{-1}$ can be obtained. Suppose next that $\phi(t)\notin[-c,c]$. Then for $\xi>0$ and $\delta>0$ small enough:
$|\dot{\phi}|>\frac{1}{2}|\sin c|\rho_{1}\geq\frac{1}{2}|\sin c|\xi^{-1}\epsilon>0,$
from which we estimate $T\leq C\epsilon^{-1}$. In combination, these estimates prove that $r_{1}(T)\in\partial\mathcal{A}^{+}(\epsilon)$ with $0<T(r_{1}(0),\epsilon)\leq C\epsilon^{-1}$, $C>0$ large enough, for all $0<\epsilon\ll 1$. ∎
Figure 6. Illustration of the phase portrait of (4.25) on the region $\mathcal{A}^{+}(\epsilon)$ (orange). Here $r_{1}\in\epsilon\mathcal{X}_{2}^{+}$ (blue) is obtained from $x_{2}\in\mathcal{X}_{2}^{+}$ in Proposition 3.3, cf. the change of coordinates in (4.4). We fix $\mathcal{X}_{2}^{+}$ so large so that $\epsilon\mathcal{X}_{2}^{+}$ has a non-empty intersection with $\mathcal{A}^{+}(\epsilon)$ for all $0<\epsilon\ll 1$. Then we extend the unstable manifold of $E^{+}(\epsilon)$ in Proposition 3.3 as a graph over $\mathcal{A}^{+}(\epsilon)$ by using the forward flow of (4.25) (with initial conditions in $\epsilon\mathcal{X}_{2}^{+}$), see Proposition 4.8.
Now, let $r_{1}(t)\in\mathcal{A}^{+}(\epsilon)$, $t\in[0,T]$, denote a solution of (4.25) with $r_{1}(0)\in\mathcal{A}^{+}(\epsilon)$. Here $T=T(r_{1}(0),\epsilon)>0$ is as in Lemma 4.6. Then we denote by $t\mapsto\widetilde{\mathcal{M}}_{1}^{N}(t,r_{1}(0),\epsilon)$, $t\in[0,T]$, the fundamental matrix associated with the linear problem
$\dot{\widetilde{\zeta}}_{1}(t)$ $=\left(\Omega+r_{1}(t)\widetilde{W}_{11}^{\pm,N}(r_{1}(t),\epsilon r_{1}(t)^{-1})\right)\widetilde{\zeta}_{1}(t),\quad t\in[0,T],$ (4.29)
with $\widetilde{\mathcal{M}}_{1}^{N}(0,r_{1}(0),\epsilon)=\operatorname{Id}$. Here we have used that $\epsilon_{1}=\epsilon r_{1}^{-1}$. In the following, we fix $\xi>0,\delta>0$ small enough.
Lemma 4.7.
The fundamental matrix $\widetilde{\mathcal{M}}_{1}^{N}(t,r_{1}(0),\epsilon)$ has the following expansion
$\widetilde{\mathcal{M}}_{1}^{N}(t,r_{1}(0),\epsilon)=\left(\frac{r_{1}(0)^{2}-\epsilon^{2}}{r_{1}(t)^{2}-\epsilon^{2}}\right)^{\frac{2+b}{2}}\exp(\Omega t)\mathcal{O}(1),\quad t\in[0,T].$ (4.30)
In particular,
$\|\widetilde{\mathcal{M}}_{1}^{N}(t,r_{1}(0),\epsilon)\widetilde{\mathcal{M}}_{1}^{N}(s,r_{1}(0),\epsilon)^{-1}\|\leq C\epsilon^{-2-b},$ (4.31)
for some (new) constant $C>0$, for all $0<\epsilon\ll 1,\,0\leq s\leq t\leq T$.
Proof.
We write
$\widetilde{\mathcal{M}}_{1}^{N}(t,r_{1}(0),\epsilon)=\left(\frac{r_{1}(0)^{2}-\epsilon^{2}}{r_{1}(t)^{2}-\epsilon^{2}}\right)^{\frac{2+b}{2}}{\mathrm{e}}^{(\Omega+\sigma\epsilon\operatorname{Id})t}\widehat{\mathcal{M}}_{1}^{N}(t,r_{1}(0),\epsilon).$ (4.32)
A simple calculation then shows that $\widehat{\mathcal{M}}_{1}^{N}(t,r_{1}(0),\epsilon)$ is the fundamental matrix associated with the linear problem
$\dot{\widehat{\zeta}}_{1}(t)$ $=\mathcal{O}(r_{1}(t)^{2})\widehat{\zeta}_{1},$
satisfying $\widehat{\mathcal{M}}_{1}^{N}(0,r_{1}(0),\epsilon)=\operatorname{Id}$. Here we have used (4.23), $\xi>0$ and $\delta>0$ small enough, and that $|r_{1}(t)|\in[\xi^{-1}\epsilon,\delta]$, $t\in[0,T]$, by assumption. But then
$\frac{d\widehat{\zeta}_{1}}{dr_{1}}=\mathcal{O}(1)\widehat{\zeta}_{1},$
by the chain rule, recall (4.25). It follows that $\|\widehat{\mathcal{M}}_{1}^{N}(t,r_{1}(0),\epsilon)\widehat{\mathcal{M}}_{1}^{N}(s,r_{1}(0),\epsilon)^{-1}\|$, $0\leq s\leq t\leq T$, is uniformly bounded. This proves (4.30) and (4.31), upon using (4.28) and $\|\exp({\Omega t})\|\equiv 1$. ∎
4.3. Extension of the unstable manifold
By taking the compact set $\mathcal{X}_{2}^{+}$ in Proposition 3.3 large enough, we can by the change of coordinates (4.4), write the unstable manifold of $E^{+}(\epsilon)$ in the graph form
$(y_{1},z_{1})=m_{1}^{+}(r_{1},\epsilon):=(\epsilon r_{1}^{-1})^{2}m_{2}^{+}(\epsilon^{-1}r_{1},\epsilon),$ (4.33)
on the boundary $r_{1}\in\partial^{1}\mathcal{A}^{+}(\epsilon)$ of $\mathcal{A}^{+}(\epsilon)$ defined by $|r_{1}|=\xi^{-1}\epsilon$. We then use the flow with (4.33) as initial conditions to extend the graph (4.33) over the region $r_{1}\in\mathcal{A}^{+}(\epsilon)$.
Notice that since $m_{2}^{+}$ is asymptotic to the series (3.9), and the series (4.10) coincides with (3.9) upon the change of coordinates (4.4), it follows that (4.33) becomes
$\widetilde{\zeta}_{1}(r_{1},\epsilon)=\epsilon^{N}\widetilde{m}_{1}^{+,N}(r_{1},\epsilon),$ (4.34)
with respect to the $\widetilde{\zeta}_{1}$-coordinate, for $r_{1}\in\partial^{1}\mathcal{A}^{+}(\epsilon)$ and some locally defined real-analytic function $\widetilde{m}_{1}^{+,N}(\cdot,\epsilon)$, recall (4.19). We then use Lemma 4.7 to write (4.21) as a fixed-point equation:
$\widetilde{\zeta}_{1}(t)$ $=\widetilde{\mathcal{M}}_{1}^{N}(t,r_{1}(0),\epsilon)\epsilon^{N}\widetilde{m}_{1}^{+,N}(r_{1}(0),\epsilon)$ (4.35)
$+\int_{0}^{t}\widetilde{\mathcal{M}}_{1}^{N}(t,r_{1}(0),\epsilon)\widetilde{\mathcal{M}}_{1}^{N}(s,r_{1}(0),\epsilon)^{-1}\epsilon^{N}\widetilde{W}_{10}^{\pm,N}(r_{1}(s),\widetilde{\zeta}_{1}(s),\epsilon r_{1}(s)^{-1})ds,$
where $r_{1}(t)\in\mathcal{A}^{+}(\epsilon)$ is the solution of (4.25) in $\mathcal{A}^{+}(\epsilon)$ with $r_{1}(0)\in\partial^{1}\mathcal{A}^{+}(\epsilon)$. In fact, to ensure that $r_{1}(t)\in\mathcal{A}^{+}(\epsilon)$ for $t\in[0,T]$, we restrict attention to initial conditions $r_{1}(0)=\xi^{-1}\epsilon e^{i\phi}$ with $\phi\in(\phi_{-},\phi_{+})$, recall (4.27). See also Fig. 6. In this way, we parameterize a subset $\widetilde{\mathcal{A}}(\epsilon)\subset\mathcal{A}^{+}(\epsilon)$ through $t\in[0,T]$ and $\phi\in(\phi_{-},\phi_{+})$. The fact that $\widetilde{\mathcal{A}}(\epsilon)\neq\mathcal{A}^{+}(\epsilon)$ is not important. Indeed, $\widetilde{\mathcal{A}}(\epsilon)$ clearly contains a region (similar to $\mathcal{A}^{+}(\epsilon)$, see Fig. 6) with $|r_{1}|\in[\widetilde{\xi}^{-1}\epsilon,\widetilde{\delta}$], $\phi\in(-\frac{\pi+\chi}{2},\frac{\pi+\chi}{2})$, where $0<\widetilde{\xi}<\xi$ and $0<\widetilde{\delta}<\delta$, for all $0<\epsilon\ll 1$, and we can just work with this region instead.
Proposition 4.8.
The unstable manifold of $E^{+}(\epsilon)$ is a graph over $r_{1}\in\mathcal{A}(\epsilon)$:
$(y_{1},z_{1})=m_{1}^{+}(r_{1},\epsilon),\quad r_{1}\in\mathcal{A}(\epsilon),$ (4.36)
for all $0<\epsilon<\epsilon_{0}$, $0<\epsilon_{0}\ll 1$. Here $m_{1}^{+}(\cdot,\epsilon):\mathcal{A}(\epsilon)\rightarrow\mathbb{C}^{2}$ is real-analytic with respect to $r_{1}\in\mathcal{A}^{+}(\epsilon)$ and $C^{\infty}$ with respect to $\epsilon\in[0,\epsilon_{0})$. In particular,
$m_{1}^{+}(r_{1},0)=\psi_{1,0}^{+}(r_{1}),\quad r_{1}\in S^{+}.$ (4.37)
Proof.
We work in the $\widetilde{\zeta}_{1}$-coordinates and consider the fixed-point formulation (4.35). It suffices to estimate the solution. Fix $C>0$ and let $T_{0}\leq T(r_{1}(0),\epsilon)$ be so that $|\widetilde{\zeta}_{1}(t)|\leq C$ for all $t\in[0,T_{0}]$. Then by (4.31), we conclude from (4.35) that
$|\widetilde{\zeta}_{1}(t)|\leq\mathcal{O}(1)\epsilon^{N-2-b}+\mathcal{O}(1)\epsilon^{N-2-b}T_{0}\leq\mathcal{O}(1)\epsilon^{N-3-b},$ (4.38)
for all $t\in[0,T_{0}]$, upon using (4.28), and that $\epsilon T$ is uniformly bounded by Lemma 4.6, see (4.28). We therefore take $N>3+b$ and $T_{0}=T$ for all $0<\epsilon\ll 1$. This gives the existence of the graph (4.36) for all $0<\epsilon<\epsilon_{0}$, $0<\epsilon_{0}\ll 1$. Notice in particular that (4.37) follows from (4.20) and (4.38), see also (4.22). Finally, for the smoothness with respect to $\epsilon$, we can proceed as in the proof of Proposition 3.3.
5. A dynamical systems approach to the difference
It now follows from Proposition 3.3 and Proposition 4.8, see also (4.3) and (4.1), that we have invariant manifolds of (1.5) as graphs
$(y,z)$ $=m^{\pm}(x,\epsilon),\quad x\in S^{\pm},$ (5.1)
respectively. (Although Proposition 4.8 only relates to the unstable manifold of $E^{+}$, we can obtain a similar result for the stable manifold of $E^{-}$ upon time-reversal.) By taking the sets $\mathcal{X}^{\pm}$ large enough, we can replace $S^{\pm}$ by their closures, $\operatorname{Cl}(S^{\pm})$, in (5.1). We then consider the difference
$\Delta m:=m^{+}(x,\epsilon)-m^{-}(x,\epsilon),\quad\Delta m=:(\Delta y,\Delta z),$
on the common domain $x\in S^{+}\cap S^{-}$, which by using the closures of $S^{\pm}$, includes a subset $i[-\delta,\delta]$, $\delta>0$ small enough, of the imaginary axis $i\mathbb{R}\subset\mathbb{C}$. Each $m^{\pm}$ is third order in the following sense
$m^{\pm}(x,\epsilon)=\mathcal{O}(|(x,\epsilon)|^{3}),$
for $(x,\epsilon)\to(0,0)$ in $S^{\pm}\times[0,\epsilon_{0})$, respectively. This follows from the analysis in the charts.
Now, $(y,z)=m^{+}(x,\epsilon)$, $x\in S^{+}$, and $(y,z)=m^{-}(x,\epsilon)$, $x\in S^{-}$, are invariant manifolds, satisfying the invariance equation
$\frac{d}{dx}\begin{pmatrix}y\\ z\end{pmatrix}=J(x,y,z,\epsilon),$
with
$J(x,y,z,\epsilon):=\frac{1}{x^{2}-\epsilon^{2}+a(y^{2}+z^{2})+F}\left(\left(\Omega+b(\sigma\epsilon-x)\operatorname{Id}\right)\begin{pmatrix}y\\ z\end{pmatrix}+\begin{pmatrix}G\\ H\end{pmatrix}\right)$,
where $W=W(x,y,z,\epsilon)$, $W=F,G,H$. In particular, we note that $J(x,m^{\pm}(x,\epsilon),\epsilon)$ are both well-defined on $x\in S^{+}\cap S^{-}$ for all $0<\epsilon\ll 1$. Consequently,
$\frac{d}{dx}\Delta m$ $=J(x,m^{-}+\Delta m,\epsilon)-J(x,m^{-},\epsilon)$ (5.2)
$=:\begin{pmatrix}A_{11}(x,\epsilon)&A_{12}(x,\epsilon)\\ A_{21}(x,\epsilon)&A_{22}(x,\epsilon)\end{pmatrix}\Delta m,$
by the mean value theorem, where
$A_{11}(x,\epsilon)$ $=\frac{b(-x+\epsilon\sigma)+G^{\prime}_{y}(x,y_{*},z_{*},\epsilon)}{x^{2}-\epsilon^{2}+a(y_{*}^{2}+z_{*}^{2})+F(x,y_{*},z_{*},\epsilon)}$
$-\frac{b(-x+\epsilon\sigma)y_{*}+z_{*}+G(x,y_{*},z_{*},\epsilon)}{(x^{2}-\epsilon^{2}+a(y_{*}^{2}+z_{*}^{2})+F(x,y_{*},z_{*},\epsilon))^{2}}(2ay_{*}+F^{\prime}_{y}(x,y_{*},z_{*},\epsilon),$
$A_{12}(x,\epsilon)$ $=\frac{1+G^{\prime}_{z}(x,y_{*},z_{*},\epsilon)}{x^{2}-\epsilon^{2}+a(y_{*}^{2}+z_{*}^{2})+F(x,y_{*},z_{*},\epsilon)}$
$-\frac{b(-x+\epsilon\sigma)y_{*}+z_{*}+G(x,y_{*},z_{*},\epsilon)}{(x^{2}-\epsilon^{2}+a(y_{*}^{2}+z_{*}^{2})+F(x,y_{*},z_{*},\epsilon))^{2}}(2az_{*}+F^{\prime}_{z}(x,y_{*},z_{*},\epsilon),$
and
$A_{21}(x,\epsilon)$ $=\frac{-1+H^{\prime}_{y}(x,y_{*},z_{*},\epsilon)}{x^{2}-\epsilon^{2}+a(y_{*}^{2}+z_{*}^{2})+F(x,y_{*},z_{*},\epsilon)}$
$-\frac{b(-x+\epsilon\sigma)z_{*}-y_{*}+H(x,y_{*},z_{*},\epsilon)}{(x^{2}-\epsilon^{2}+a(y_{*}^{2}+z_{*}^{2})+F(x,y_{*},z_{*},\epsilon))^{2}}(2ay_{*}+F^{\prime}_{y}(x,y_{*},z_{*},\epsilon),$
$A_{22}(x,\epsilon)$ $=\frac{b(-x+\epsilon\sigma)+H^{\prime}_{z}(x,y_{*},z_{*},\epsilon)}{x^{2}-\epsilon^{2}+a(y_{*}^{2}+z_{*}^{2})+F(x,y_{*},z_{*},\epsilon)}$
$-\frac{b(-x+\epsilon\sigma)z_{*}-y_{*}+H(x,y_{*},z_{*},\epsilon)}{(x^{2}-\epsilon^{2}+a(y_{*}^{2}+z_{*}^{2})+F(x,y_{*},z_{*},\epsilon))^{2}}(2az_{*}+F^{\prime}_{z}(x,y_{*},z_{*},\epsilon).$
Here
$(y_{*},z_{*})=(y_{*},z_{*})(x,\epsilon)=\mathcal{O}(|(x,\epsilon)|^{3}),$
for $(x,\epsilon)\to(0,0)$ in $(S^{+}\cap S^{-})\times[0,\epsilon_{0})$, are (four different pairs of) known real-analytic functions (the intermediate points in the application of the mean value theorem), depending smoothly on $\epsilon\in[0,\epsilon_{0})$ (in the same sense as $m^{\pm}$ in the charts above). It will not be important to distinguish the pairs, so we use the same symbols for all of them. Notice by (1.3):
$F(x,y_{*},z_{*},\epsilon)=F_{0}x^{3}+\mathcal{O}(|(x,\epsilon)|^{4}),$ (5.3)
for $(x,\epsilon)\to(0,0)$ in $(S^{+}\cap S^{-})\times[0,\epsilon_{0})$.
Now, the quantities $A_{\alpha\beta}$ are each of the form “0/00/0” for $(x,\epsilon)=(0,0)$, so we study the resulting system using blowup. It will suffice to work along the imaginary axis, so we put
$x=iv,$
and use the real blowup:
$\rho\geq 0,\,(\breve{\rho},\breve{\epsilon})\in S^{1}\mapsto\begin{cases}v=\rho\breve{v},\\ \epsilon=\rho\breve{\epsilon}.\end{cases}$ (5.4)
Remark 5.1.
Notice that in comparison with the earlier blowup, we only blowup $x$ and $\epsilon$ and we focus on $x\in i\mathbb{R}$. Therefore (5.4) is real. Notice also, that we abuse notation slightly in (5.4) by redefining $\breve{\epsilon}$ (which was used in a different (complex) context in (4.2)).
Moreover, as (1.5) is real-analytic, so that $(x(t),y(t),z(t))$, $t\in\mathbb{R}$, is a solution if and only if $(\overline{x}(t),\overline{y}(t),\overline{z}(t))$, $t\in\mathbb{R}$, is a solution, we have
${A_{\alpha\beta}(-iv,\epsilon)}=\overline{A_{\alpha\beta}}(iv,\epsilon),$ (5.5)
and in particular
$\begin{cases}\Delta y(-iv)=\overline{\Delta y}(iv),\\ \Delta z(-iv)=\overline{\Delta z}(iv),\end{cases}$
where the bar denotes complex conjugation, for all $v\in[-\delta,\delta]$. The system
$\frac{d\Delta y}{dv}$ $=i\left(A_{11}(iv,\epsilon)\Delta y+A_{12}(iv,\epsilon)\Delta z\right),$ (5.6)
$\frac{d\Delta z}{dv}$ $=i\left(A_{21}(iv,\epsilon)\Delta y+A_{22}(iv,\epsilon)\Delta z\right),$
obtained by substituting $x=iv$ into (5.2), is therefore also invariant with respect to the time-reversal symmetry
$(v,\Delta y,\Delta z)\mapsto(-v,\overline{\Delta y},\overline{\Delta z}).$ (5.7)
We work in the charts $\breve{v}=1$, $\breve{\epsilon}=1$ and $\breve{v}=-1$ with chart-specific coordinates $(\rho_{1},\epsilon_{1})$, $(\rho_{2},v_{2})$ and $(\rho_{3},\epsilon_{3})$, respectively, defined by:
$\breve{v}=1:\quad\begin{cases}v=\rho_{1},\\ \epsilon=\rho_{1}\epsilon_{1},\end{cases}$
$\breve{\epsilon}=1:\quad\begin{cases}v=\rho_{2}v_{2},\\ \epsilon=\rho_{2},\end{cases}$
$\breve{v}=-1:\quad\begin{cases}v=-\rho_{3},\\ \epsilon=\rho_{3}\epsilon_{3}.\end{cases}$
The change of coordinates are given by the expressions
$\begin{cases}\rho_{1}=\rho_{2}v_{2},\\ \epsilon_{1}=v_{2}^{-1},\end{cases}\qquad\begin{cases}\rho_{3}=\rho_{2}(-v_{2}),\\ \epsilon_{3}=(-v_{2})^{-1}.\end{cases}$ (5.8)
In the scaling chart: $\breve{\epsilon}=1$, we find
$A_{11}(\epsilon iv_{2},\epsilon)$ $=\epsilon^{-2}\left(\frac{b\epsilon(-iv_{2}+\sigma)+\mathcal{O}(\epsilon^{2})}{-v_{2}^{2}-1-\epsilon F_{0}iv_{2}^{3}+\mathcal{O}(\epsilon^{2})}-\frac{\mathcal{O}(\epsilon^{3})}{\left(-v_{2}^{2}-1-\epsilon F_{0}iv_{2}^{3}+\mathcal{O}(\epsilon^{2})\right)^{2}}\right),$ (5.9)
$A_{12}(\epsilon iv_{2},\epsilon)$ $=\epsilon^{-2}\left(\frac{1+\mathcal{O}(\epsilon^{2})}{-v_{2}^{2}-1-\epsilon F_{0}iv_{2}^{3}+\mathcal{O}(\epsilon^{2})}-\frac{\mathcal{O}(\epsilon^{3})}{\left(-v_{2}^{2}-1-\epsilon F_{0}iv_{2}^{3}+\mathcal{O}(\epsilon^{2})\right)^{2}}\right),$
$A_{21}(\epsilon iv_{2},\epsilon)$ $=\epsilon^{-2}\left(\frac{-1+\mathcal{O}(\epsilon^{2})}{-v_{2}^{2}-1-\epsilon F_{0}iv_{2}^{3}+\mathcal{O}(\epsilon^{2})}-\frac{\mathcal{O}(\epsilon^{3})}{\left(-v_{2}^{2}-1-\epsilon F_{0}iv_{2}^{3}+\mathcal{O}(\epsilon^{2})\right)^{2}}\right),$
$A_{22}(\epsilon iv_{2},\epsilon)$ $=\epsilon^{-2}\left(\frac{b\epsilon(-iv_{2}+\sigma)+\mathcal{O}(\epsilon^{2})}{-v_{2}^{2}-1-\epsilon F_{0}iv_{2}^{3}+\mathcal{O}(\epsilon^{2})}-\frac{\mathcal{O}(\epsilon^{3})}{\left(-v_{2}^{2}-1-\epsilon F_{0}iv_{2}^{3}+\mathcal{O}(\epsilon^{2})\right)^{2}}\right).$
Here we have used (5.3). This leads to the following system
$\frac{d\Delta y}{dv_{2}}$ $=\frac{i}{\epsilon(-v_{2}^{2}-1-\epsilon F_{0}iv_{2}^{3})}\left((b\epsilon(-iv_{2}+\sigma)+\mathcal{O}(\epsilon^{2}))\Delta y+(1+\mathcal{O}(\epsilon^{2}))\Delta z\right),$
$\frac{d\Delta z}{dv_{2}}$ $=\frac{i}{\epsilon(-v_{2}^{2}-1-\epsilon F_{0}iv_{2}^{3}))}\left((b\epsilon(-iv_{2}+\sigma)+\mathcal{O}(\epsilon^{2}))\Delta z-(1+\mathcal{O}(\epsilon^{2}))\Delta y\right),$
for $v_{2}\in I$, with $I\subset\mathbb{R}$ compact. To study this system, we use the following equivalent first order formulation:
$v_{2}^{\prime}$ $=-\epsilon,$ (5.10)
$\Delta y^{\prime}$ $=\frac{i}{v_{2}^{2}+1+\epsilon F_{0}iv_{2}^{3}}\left((b\epsilon(-iv_{2}+\sigma)+\mathcal{O}(\epsilon^{2}))\Delta y+(1+\mathcal{O}(\epsilon^{2}))\Delta z\right),$
$\Delta z^{\prime}$ $=\frac{i}{v_{2}^{2}+1+\epsilon F_{0}iv_{2}^{3}}\left((b\epsilon(-iv_{2}+\sigma)+\mathcal{O}(\epsilon^{2}))\Delta z-(1+\mathcal{O}(\epsilon^{2}))\Delta y\right).$
The advantage of working with (5.10) is that we can use dynamical systems technique. In particular, $\epsilon=0$ is well-defined:
$v_{2}^{\prime}$ $=0,$
$\Delta y^{\prime}$ $=\frac{i}{v_{2}^{2}+1}\Delta z,$
$\Delta z^{\prime}$ $=-\frac{i}{v_{2}^{2}+1}\Delta y.$
We notice that $(v_{2},0,0)$ is a normally hyperbolic line, the linearization having eigenvalues $0,\pm 1$. (In this sense, the imaginary axis is a hyperbolic path, see [26, 32].) In particular, the two-dimensional sets given by
$(v_{2},-i\Delta z,\Delta z)\quad\mbox{and}\quad(v_{2},i\Delta z,\Delta z),$ (5.11)
define invariant stable and unstable manifolds, respectively, of the line $(v_{2},0,0)$, $v_{2}\in I$, for $\epsilon=0$. For $0\leq\epsilon\ll 1$, we therefore have stable and unstable manifolds (which are line bundles by the linearity):
$\begin{cases}\Delta y=Y_{2}^{s}(v_{2},\epsilon)\Delta z,\\ \Delta y=Y_{2}^{u}(v_{2},\epsilon)\Delta z,\end{cases}$ (5.12)
where
$\begin{cases}Y_{2}^{s}(v_{2},\epsilon)=-i+\mathcal{O}(\epsilon),\\ Y_{2}^{u}(v_{2},\epsilon)=i+\mathcal{O}(\epsilon).\end{cases}$
are $C^{n}$-smooth (for any $n\in\mathbb{N}$) with respect to $v_{2}\in I$, $0\leq\epsilon<\epsilon_{0}(n)$, $0<\epsilon_{0}(n)\ll 1$, with $I\subset\mathbb{R}$ a fixed compact set. This follows from Fenichel’s theory, see e.g. [28]. In fact, by the symmetry (5.7), (5.10) has a time-reversal symmetry $(v_{2},\Delta y,\Delta z,t)\mapsto(-v_{2},\overline{\Delta y},\overline{\Delta z},-t)$, and we can therefore take
$Y_{2}^{u}(v_{2},\epsilon):=\overline{Y_{2}^{s}}(-v_{2},\epsilon).$
In turn,
$\begin{pmatrix}\Delta y\\ \Delta z\end{pmatrix}$ $=\begin{pmatrix}Y_{2}^{s}(v_{2},\epsilon)&\overline{Y_{2}^{s}}(-v_{2},\epsilon)\\ 1&1\end{pmatrix}\begin{pmatrix}\Delta s\\ \Delta u\end{pmatrix}$, (5.13)
defines a $C^{n}$-smooth $\epsilon$-dependent change of coordinates $(v_{2},\Delta y,\Delta z)\mapsto(v_{2},\Delta s,\Delta u)$, that diagonalizes (5.10):
$v_{2}^{\prime}$ $=-\epsilon,$ (5.14)
$\Delta s^{\prime}$ $=\frac{1}{v_{2}^{2}+1}\left(-1+\mathcal{O}(\epsilon)\right)\Delta s,$
$\Delta u^{\prime}$ $=\frac{1}{v_{2}^{2}+1}\left(1+\mathcal{O}(\epsilon)\right)\Delta u.$
In these coordinates, the time-reversal symmetry takes the form
$(v_{2},\Delta s,\Delta u,t)\mapsto(-v_{2},\overline{\Delta u},\overline{\Delta s},-t).$
The system (5.14) can be viewed as a Fenichel normal form, see [28, Section 3]. The following result shows that the diagonalization (5.13) can be “globalized” beyond a compact subset $v_{2}\in I$ (by working in the directional charts $\breve{v}=\pm 1$ associated with the blowup (5.4)).
Proposition 5.2.
Fix any $n\in\mathbb{N}$ and consider (5.2) with $x=iv$, $v\in[-\delta,\delta]$, $\delta>0$ small enough. Then there exists a locally defined $\epsilon$-dependent and $v$-fibered change of coordinates $(v,\Delta y,\Delta z)\mapsto(v,\Delta s,\Delta u)$ defined by
$\begin{pmatrix}\Delta y\\ \Delta z\end{pmatrix}=\begin{pmatrix}-i+\mathcal{O}(\rho^{2})&i+\mathcal{O}(\rho^{2})\\ 1&1\end{pmatrix}\begin{pmatrix}\Delta s\\ \Delta u\end{pmatrix}$,
that diagonalizes (5.2):
$\frac{d\Delta s}{dv}$ $=\frac{1}{-v^{2}-\epsilon^{2}}\left(-1+\frac{F_{0}iv^{3}}{v^{2}+\epsilon^{2}}+bv+i\epsilon b\sigma+\mathcal{O}(\rho^{2})\right)\Delta s,$ (5.15)
$\frac{d\Delta u}{dv}$ $=\frac{1}{-v^{2}-\epsilon^{2}}\left(1-\frac{F_{0}iv^{3}}{v^{2}+\epsilon^{2}}+bv+i\epsilon b\sigma+\mathcal{O}(\rho^{2})\right)\Delta u.$
Here all $\mathcal{O}(\rho^{2})$-terms are $C^{n}$-smooth with respect to $\rho\in[0,\delta]$, $(\breve{v},\breve{\epsilon})\in S^{1}$ (i.e. they are smooth functions of $(\rho_{1},\epsilon_{1})$, $(v_{2},\rho_{2}=\epsilon)$ and $(\rho_{3},\epsilon_{3})$ in the charts). Moreover, the system (5.15) is invariant with respect to the time-reversal symmetry $(v,\Delta s,\Delta u)\mapsto(-v,\overline{\Delta u},\overline{\Delta s})$ for all $0<\epsilon\ll 1$.
Proof.
In order to prove the statement, we will extend (globalize) the stable and unstable manifolds of (5.10), recall (5.12), beyond the scaling chart ($\breve{\epsilon}=1$). For this purpose, it is convenient to work in projective coordinates: For $\Delta z\neq 0$, define
$Y:=\frac{\Delta y}{\Delta z}.$
Then a simple calculation shows that
$\frac{dY}{dv}=i\left(A_{12}(iv,\epsilon)+(A_{11}(iv,\epsilon)-A_{22}(iv,\epsilon))Y-A_{21}(iv,\epsilon)Y^{2}\right),$
which is invariant with respect to $(v,Y)\mapsto(-v,\overline{Y})$, recall (5.5). In the $\breve{\epsilon}=1$-chart, we obtain the first order system
$v_{2}^{\prime}$ $=-\epsilon,$ (5.16)
$Y^{\prime}$ $=\frac{i}{v_{2}^{2}+1+\epsilon F_{0}iv_{2}^{3}}\left(1+\mathcal{O}(\epsilon^{2})+\mathcal{O}(\epsilon^{2})Y+(1+\mathcal{O}(\epsilon^{2}))Y^{2}\right),$
$\epsilon^{\prime}$ $=0,$
by (5.9). Now, for $\epsilon=0$ we have critical manifolds defined by $(v_{2},\mp i,0)$ which are normally hyperbolic. Indeed, the linearization about any point $(v_{2},\mp i,0)$ has a single nontrivial eigenvalue of the form
$\frac{i}{v_{2}^{2}+1}2(\mp i)=\frac{\pm 2}{v_{2}^{2}+1},$ (5.17)
respectively. We focus on Y=iY=-i (and use the symmetry to study $Y=i$). By (5.11), Y=iY=-i corresponds to the stable manifold in the $(v_{2},\Delta y,\Delta z)$-space for $\epsilon=0$ and notice that it is unstable for (5.16) by (5.17). Now, by working in a compact set $v_{2}\in I$, we have a repelling slow manifold:
$Y=-i+\mathcal{O}(\epsilon^{2}),$ (5.18)
as a $C^{n}$-smooth graph over $(v_{2},\epsilon)\in I\times[0,\epsilon_{0})$, $0<\epsilon_{0}\ll 1$. Next, we consider the coordinates of the $\breve{v}=1$-chart:
$A_{11}(i\rho_{1},\rho_{1}\epsilon_{1})$ $=\frac{1}{\rho_{1}^{2}\left[-1-\epsilon_{1}^{2}-F_{0}i\rho_{1}\right]}\left(b\rho_{1}(-i+\epsilon_{1}\sigma)+\mathcal{O}(\rho_{1}^{2})\right),$
$A_{12}(i\rho_{1},\rho_{1}\epsilon_{1})$ $=\frac{1}{\rho_{1}^{2}\left[-1-\epsilon_{1}^{2}-F_{0}i\rho_{1}\right]}\left(1+\mathcal{O}(\rho_{1}^{2})\right),$
$A_{21}(i\rho_{1},\rho_{1}\epsilon_{1})$ $=\frac{1}{\rho_{1}^{2}\left[-1-\epsilon_{1}^{2}-F_{0}i\rho_{1}\right]}\left(-1+\mathcal{O}(\rho_{1}^{2})\right),$
$A_{22}(i\rho_{1},\rho_{1}\epsilon_{1})$ $=\frac{1}{\rho_{1}^{2}\left[-1-\epsilon_{1}^{2}-F_{0}i\rho_{1}\right]}\left(b\rho_{1}(-i+\epsilon_{1}\sigma)+\mathcal{O}(\rho_{1}^{2})\right).$
Consequently,
$\frac{dY}{d\rho_{1}}$ $=\frac{i}{\rho_{1}^{2}\left[-1-\epsilon_{1}^{2}-F_{0}i\rho_{1}\right]}\left(1+\mathcal{O}(\rho_{1}^{2})+\mathcal{O}(\rho_{1}^{2})Y+(1+\mathcal{O}(\rho_{1}))Y^{2}\right),$
or equivalently
$\dot{\rho}_{1}$ $=-\rho_{1}^{2},$ (5.19)
$\dot{Y}$ $=\frac{i}{\left[1+\epsilon_{1}^{2}+F_{0}i\rho_{1}\right]}\left(1+\mathcal{O}(\rho_{1}^{2})+\mathcal{O}(\rho_{1}^{2})Y+(1+\mathcal{O}(\rho_{1}))Y^{2}\right),$
$\dot{\epsilon}_{1}$ $=\rho_{1}\epsilon_{1}.$
In comparison with (5.16), we have used a desingularization of time corresponding to division of the right hand side by $\epsilon_{1}^{2}$, i.e. $\dot{(\cdot)}=\epsilon_{1}^{-2}(\cdot)^{\prime}$. For (5.19), we have partially hyperbolic singularities $(0,\pm i,0)$, that are stable and unstable, respectively. Indeed the linearization around $(0,\pm i,0)$ have eigenvalues $0,\mp 2,0$. We focus on (0,i,0)(0,-i,0), having a center manifold:
$Y=-i+\mathcal{O}(\rho_{1}^{2}),$ (5.20)
as a $C^{n}$-smooth graph over $\rho_{1}\in[0,\delta],\epsilon_{1}\in[0,\xi]$, with $\delta>0$ and $\xi>0$ small enough. The center manifold, which is repelling for (5.19), is foliated by $\epsilon=\rho_{1}\epsilon_{1}$ constant.
The results in the $\breve{v}=-1$-chart are similar. Here we also obtain a partially hyperbolic equilibrium point $(\rho_{3},Y,\epsilon_{3})=(0,-i,0)$ and a two-dimensional center manifold:
$Y=-i+\mathcal{O}(\rho_{3}^{2}),$ (5.21)
as a $C^{n}$-smooth graph over $\rho_{3}\in[0,\delta],\epsilon_{3}\in[0,\xi]$, with $\delta>0$ and $\xi>0$ small enough. The center manifold is also repelling and foliated by $\epsilon=\rho_{3}\epsilon_{3}$ constant.
Now, notice that with respect to the backward flow, the center manifolds in the $\breve{v}=\pm 1$-charts and the slow manifold in the $\breve{\epsilon}=1$-chart all become attracting and moreover $v$ (in particular $v_{2}$) becomes an increasing function. Therefore, following the pioneering work of Dumortier and Roussarie [18], see also [33] by Krupa and Szmolyan, we can use the backward flow to extend a fixed center manifold in the $\breve{v}=-1$-chart, see (5.21), into the scaling chart $\breve{\epsilon}=1$ (since $\rho_{3}=-p$ is decreasing) and continue this as a smooth invariant manifold there upon using the slow manifold, see (5.18), which is also attracting for the backward flow. Moreover, from here we can extend the resulting manifold into the $\breve{v}=1$-chart by using the attracting center manifold, see (5.20). In this way, we obtain a global invariant manifold
$Y=Y^{s}(v,\epsilon)=-i+\mathcal{O}(\rho^{2}),$
with $\mathcal{O}(\rho^{2})$ being $C^{n}$-smooth on the cylinder (i.e. in the charts).
Next, we obtain the unstable manifold by applying the time-reversal symmetry:
$Y=Y^{u}(v,\epsilon):=\overline{Y^{s}}(-v,\epsilon).$
Then the desired diagonalization is given by
$\begin{pmatrix}\Delta y\\ \Delta z\end{pmatrix}=\begin{pmatrix}Y^{s}(v,\epsilon)&\overline{Y^{s}}(-v,\epsilon)\\ 1&1\end{pmatrix}\begin{pmatrix}\Delta s\\ \Delta u\end{pmatrix}$ (5.22)
which brings the stable and unstable manifolds to $\Delta u=0$ and $\Delta s=0$, respectively. In turn, it follows that the linear system has been diagonalized, so that $\frac{d\Delta}{dv}$ only depends upon $\Delta$ for $\Delta=\Delta s,\Delta u$. Therefore to compute ${d\Delta s}/{dv}$ we can just put $\Delta u=0$ so that $\Delta z=\Delta s$ and set $\Delta y=Y^{s}(v,\epsilon)\Delta z$ (cf. (5.22)) into the $\Delta z$-equation:
$\frac{d\Delta s}{dv}$ $=i\left(A_{21}(iv,\epsilon)Y^{s}(v,\epsilon)+A_{22}(iv,\epsilon)\right)\Delta s.$
${d\Delta u}/{dv}$ is obtained in the same way:
$\frac{d\Delta u}{dv}$ $=i\left(A_{21}(iv,\epsilon)\overline{Y^{s}}(-v,\epsilon)+A_{22}(iv,\epsilon)\right)\Delta u.$
Now, using the expansions of $A_{21},A_{22}$ and $Y^{s}$ in the charts, we obtain the desired result. In particular, in the $\breve{v}=1$-chart, we find that
$\frac{d\Delta s}{d\rho_{1}}$ $=\frac{1}{\rho_{1}^{2}\left[-1-\epsilon_{1}^{2}-F_{0}i\rho_{1}\right]}\left(-1+ib\rho_{1}(-i+\epsilon_{1}\sigma)+\mathcal{O}(\rho_{1}^{2})\right)\Delta s,$
$=\frac{1}{\rho_{1}^{2}\left[-1-\epsilon_{1}^{2}\right]}\left(-1+\frac{F_{0}i\rho_{1}}{1+\epsilon_{1}^{2}}+b\rho_{1}+i\rho_{1}\epsilon_{1}b\sigma+\mathcal{O}(\rho_{1}^{2})\right)\Delta s,$
using the expansion
$\left[1+\epsilon_{1}^{2}+F_{0}i\rho_{1}\right]^{-1}=\frac{1}{1+\epsilon_{1}^{2}}-\frac{F_{0}i\rho_{1}}{(1+\epsilon_{1}^{2})^{2}}+\mathcal{O}(\rho_{1}^{2}),$
uniformly on a compact domain $\epsilon_{1}\in[0,\xi]$, $\xi>0$, and similarly
$\frac{d\Delta u}{d\rho_{1}}$ $=\frac{1}{\rho_{1}^{2}\left[-1-\epsilon_{1}^{2}\right]}\left(1-\frac{F_{0}i\rho_{1}}{1+\epsilon_{1}^{2}}+b\rho_{1}+i\rho_{1}\epsilon_{1}b\sigma+\mathcal{O}(\rho_{1}^{2})\right)\Delta u.$
Figure 7. Schematic illustration of the use of blowup to perform the extension of the invariant manifolds used to diagonalize (5.2), see Proposition 5.2. There are slow manifolds (light blue) on the side of the cylinder (obtained by working in a large but fixed compact sets of the $\breve{\epsilon}=1$-chart ) and center manifolds (darker blue) on the top and bottom (working in small compact sets of the charts $\breve{v}=\pm 1$, respectively). In particular, we obtain the global invariant manifold (in blue) by fixing a center manifold on the bottom and then extend this manifold up the cylinder by using the backward flow (where the manifolds become attracting). The invariant manifolds near Y=iY=-i (blue) and $Y=i$ (red) are related by symmetry $(v,Y)\mapsto(-v,\overline{Y})$.
5.1. Computing the difference
We will now determine a formula for the difference $(\Delta y,\Delta z)$ at $x=0$ as $\epsilon\rightarrow 0$ (recall (1.13)). For this purpose, we will use (5.15) and the following result regarding the unperturbed manifolds:
Lemma 5.3.
Suppose that the unperturbed manifolds $(y,z)=\psi^{\pm}_{0}(x)$, $x\in S^{\pm}$, recall Lemma 2.1, are non-analytic at the origin (i.e. the series (2.3) is divergent for any $x\neq 0$). Then
$\psi^{+}_{0}(iv)\neq\psi_{0}^{-}(iv)\quad\forall\,v\neq 0.$ (5.23)
Proof.
The proof is by contradiction. First, we note that $\psi^{+}_{0}$ and $\psi^{-}_{0}$ satisfy the same differential equation on $x\in S^{+}\cap S^{-}$, recall Lemma 2.1. This equation is clearly regular away from $x\neq 0$. Therefore if $\psi^{+}_{0}(x)=\psi^{-}_{0}(x)$ for some $x\in S^{+}\cap S^{-}$, $x\neq 0$, then by uniqueness of solutions, we have $\psi^{+}_{0}(x)=\psi^{-}_{0}(x)$ for all $x\in B_{\xi}$, i.e. in a full neighborhood of the origin. But then by Riemann’s theorem of removable singularities, we conclude that the unperturbed manifolds are analytic at the origin.
Now, we fix $n\in\mathbb{N}$ and $\delta>0$ small enough, and suppose that the unperturbed manifolds are non-analytic at the origin. It then follows that
$\Theta(\epsilon):$ $=\begin{pmatrix}Y^{s}(\delta,\epsilon)&\overline{Y^{s}}(-\delta,\epsilon)\\ 1&1\end{pmatrix}^{-1}\left(m^{+}(i\delta,\epsilon)-m^{-}(i\delta,\epsilon)\right),$
is $C^{n}$-smooth with respect to $\epsilon\in[0,\epsilon_{0})$, and importantly:
$\Theta(0)\neq(0,0)\in\mathbb{C}^{2},$ (5.24)
by (4.37) and (5.23). Here $Y^{s}=Y^{s}(v,\epsilon)$, $v\in[-\delta,\delta]$, $\epsilon\in[0,\epsilon_{0})$, is the $C^{n}$-smooth function from Proposition 5.2.
Let $(\Theta_{s},\Theta_{u})$ denote the components of $\Theta\in\mathbb{C}^{2}$. By (5.24), it follows that either $\Theta_{s}(0)\neq 0$ or $\Theta_{u}(0)\neq 0$. Suppose the latter and consider the solution $\Delta s=\Delta s(v)$ of the $\Delta s$-equation:
$\frac{d\Delta s}{dv}=\frac{1}{-v^{2}-\epsilon^{2}}\left(-1+\frac{F_{0}iv^{3}}{v^{2}+\epsilon^{2}}+bv+i\epsilon b\sigma+\mathcal{O}(\rho^{2})\right)\Delta s,$ (5.25)
with $\Delta s(-\delta)=\overline{\Theta_{u}}$ as the initial condition. Since the system is linear, we can integrate the equation directly and if $\Theta_{u}(0)\neq 0$ then we easily conclude that $\Delta s(0)$ is exponentially large with respect to $\epsilon\rightarrow 0$ (see also Section 5.2 below). But this contradicts Proposition 3.3 and the boundedness of $m_{2}^{\pm}(0,\epsilon)$. We therefore conclude that $\Theta_{s}(0)\neq 0$ and instead use $\Delta s(\delta)=\Theta_{s}$ as an initial condition for (5.25). We obtain
$\Delta s(0)={\mathrm{e}}^{\int_{\delta}^{0}\frac{1}{-v^{2}-\epsilon^{2}}\left(-1+\frac{F_{0}iv^{3}}{v^{2}+\epsilon^{2}}+bv+i\epsilon b\sigma+\mathcal{O}(\rho^{2})\right)dv}\Theta_{s}.$ (5.26)
To determine $\Delta u(0)$ we use the symmetry: $\Delta u(v)=\overline{\Delta s}(-v)$ for all $v\in[-\delta,0]$. In particular, $\Delta u(0)=\overline{\Delta s}(0)$. We are now ready to complete the proof of the main result, Theorem 1.2.
5.2. Completing the proof of Theorem 1.2
We start by analyzing the integral in (5.26):
$\int_{\delta}^{0}\frac{1}{-v^{2}-\epsilon^{2}}\left(-1+\frac{F_{0}iv^{3}}{v^{2}+\epsilon^{2}}+bv+i\epsilon b\sigma+\mathcal{O}(\rho^{2})\right)dv.$
We expand the integrand and consider the separate terms successively. First, we directly obtain that
$\int_{\delta}^{0}\frac{1}{v^{2}+\epsilon^{2}}dv=-\epsilon^{-1}\arctan\frac{\delta}{\epsilon}=-\frac{\pi}{2}\epsilon^{-1}+\mathcal{O}(1),$ (5.27)
since
$\arctan s=\frac{\pi}{2}+\mathcal{O}(s^{-1})\quad\mbox{for $s\rightarrow\infty$};$
recall that $\delta>0$ is fixed. Next, we similarly have that
$\int_{\delta}^{0}\frac{-F_{0}iv^{3}}{(v^{2}+\epsilon^{2})^{2}}dv$ $=-\frac{1}{2}F_{0}i\left[\log(v^{2}+\epsilon^{2})+\frac{\epsilon^{2}}{\epsilon^{2}+p^{2}}\right]_{v=\delta}^{0}$
$=-\frac{1}{2}F_{0}i\left(\log\frac{\epsilon^{2}}{\delta^{2}+\epsilon^{2}}+1-\frac{\epsilon^{2}}{\epsilon^{2}+\delta^{2}}\right)$
$=F_{0}i\log\epsilon^{-1}+\mathcal{O}(1),$
and
$\int_{\delta}^{0}\frac{-bv}{v^{2}+\epsilon^{2}}dv$ $=-\frac{1}{2}b\left[\log\frac{\epsilon^{2}}{v^{2}+\epsilon^{2}}\right]_{v=\delta}^{0}$ (5.28)
$=b\log\epsilon^{-1}+\mathcal{O}(1),$
$\int_{\delta}^{0}\frac{-i\epsilon b\sigma}{v^{2}+\epsilon^{2}}dv$ $=\frac{i\pi b\sigma}{2}+\mathcal{O}(\epsilon),$ (5.29)
using (5.27) in the last integral. It remains to consider the integral
$\int_{\delta}^{0}\frac{\mathcal{O}(\rho^{2})}{-v^{2}-\epsilon^{2}}dv,$
which we analyze by working in the coordinates $(\rho_{1},\epsilon_{1})$ and $(v_{2},\epsilon)$ associated with the blowup (5.4). In particular, we split the integral
$\int_{\delta}^{0}\frac{\mathcal{O}(\rho^{2})}{-v^{2}-\epsilon^{2}}dv$ $=\left(\int_{\delta}^{\xi^{-1}\epsilon}+\int_{\xi^{-1}\epsilon}^{0}\right)\frac{\mathcal{O}(\rho^{2})}{-v^{2}-\epsilon^{2}}dv$
$=\int_{\delta}^{\xi^{-1}\epsilon}\frac{\mathcal{O}(1)}{1+(\epsilon\rho_{1}^{-1})^{2}}d\rho_{1}+\int_{\xi^{-1}}^{0}\frac{\mathcal{O}(\epsilon)}{-v_{2}^{2}-1}dv_{2}$
$=\mathcal{O}(1),$
with $\xi>0$ fixed small enough, using the substitutions $v=\rho_{1}$ and $v=\epsilon v_{2}$ in the second equality. Here we have also used that under these substitutions $\mathcal{O}(\rho^{2})$ becomes $\mathcal{O}(\rho_{1}^{2})$ and $\mathcal{O}(\epsilon^{2})$, respectively. It is easy to see that all $\mathcal{O}(1)$ terms are $C^{n}$-smooth with respect to $\epsilon$. In conclusion, we have
$\Delta s(0)$ $={\mathrm{e}}^{-\frac{\pi}{2}\epsilon^{-1}+F_{0}i\log\epsilon^{-1}+b\log\epsilon^{-1}+\mathcal{O}(1)}$
$=\epsilon^{-b}{\mathrm{e}}^{-\frac{\pi}{2}\epsilon^{-1}+F_{0}i\log\epsilon^{-1}+\mathcal{O}(1)},$
where $\mathcal{O}(1)\in\mathbb{C}$ is $C^{n}$-smooth with respect to $\epsilon$. Here we have used that $\Theta_{s}(0)\neq 0$. Then by (5.22), we have
$\begin{pmatrix}\Delta y\\ \Delta z\end{pmatrix}(0)$ $=\begin{pmatrix}Y^{s}(0,\epsilon)&\overline{Y^{s}}(-0,\epsilon)\\ 1&1\end{pmatrix}\begin{pmatrix}\Delta s(0)\\ \overline{\Delta s}(0)\end{pmatrix}$
$=2\operatorname{Re}\left(\begin{pmatrix}Y^{s}(0,\epsilon)\\ 1\end{pmatrix}\Delta s(0)\right),$
with $Y^{s}(0,\epsilon)=-i+\mathcal{O}(\epsilon)$, recall (5.18). A simple computation then shows (1.13). Notice, from the construction above, that the functions $\phi_{\alpha}$, $\alpha\in\{0,1\}$, in (1.13) are apriori only $C^{n}$. However, the manifolds are unique and smooth for $\epsilon>0$ (in particular, they are independent of our choice of $n$ and $\delta>0$). Therefore by the arbitrariness of $n\in\mathbb{N}$, we conclude that $\phi_{\alpha}$, $\alpha\in\{0,1\}$, are indeed $C^{\infty}$ with respect $\epsilon\in[0,\epsilon_{0})$.
Acknowledgement
The author is grateful to Inmaculada Baldomá and Tere M. Seara at the Polytechnic University of Catalonia, Barcelona, for the hospitality shown during two research visits in late 2024 and summer 2025. I am thankful for their kindness, interest and guidance. The author would also like to thank his close collaborators, Sam Jelbart and Peter Szmolyan, for the discussions we had during the early stages of this project in the early summer of 2024. Finally, the author thanks Sam Jelbart and Johan M. Christensen for feedback on an earlier version. The author was funded by Danish Research Council (DFF) grant 4283-00014B.
Appendix A Derivation of the system (1.1)
In [1], the authors present the following co-dimension two unfolding of the real-analytic zero-Hopf bifurcation:
$x^{\prime}$ $=x^{2}-\mu+\gamma_{3}\mu^{2}+\gamma_{4}\nu^{2}+\gamma_{5}\mu\nu+\gamma_{2}(y^{2}+z^{2})+F(x,y,z,\mu,\nu),$ (A.1)
$y^{\prime}$ $=(\nu-\beta_{1}x)y+z(\alpha_{0}+\alpha_{1}\nu+\alpha_{2}\mu+\alpha_{3}x)+G(x,y,z,\mu,\nu),$
$z^{\prime}$ $=(\nu-\beta_{1}x)z-y(\alpha_{0}+\alpha_{1}\nu+\alpha_{2}\mu+\alpha_{3}x)+H(x,y,z,\mu,\nu),$
see [1, Eq. (6)] with $(\overline{x},\overline{y},\overline{z})=:(y,z,x)$, where $\alpha_{0}\neq 0$. Here $\mu$ and $\nu$ are the local unfolding parameters, with $(\mu,\nu)=(0,0)$ corresponding to the zero-Hopf bifurcation, the linearization having eigenvalues 0, $\pm i\alpha_{0}^{2}$, and $F$, $G$, $H$ are each real-analytic and cubic with respect $(x,y,z,\mu,\nu)$. Since $\alpha_{0}\neq 0$, we can divide the right hand side by
$\alpha_{0}+\alpha_{1}\nu+\alpha_{2}\mu+\alpha_{3}x\neq 0$
for all $x,\mu,\nu$ sufficiently close to 0. This corresponds to a regular transformation of time. Then through elementary transformations of $(x,\mu,\nu)$, we can bring the system into the following form
$x^{\prime}$ $=x^{2}-\mu+\gamma_{2}(y^{2}+z^{2})+F(x,y,z,\mu,\nu),$ (A.2)
$y^{\prime}$ $=(\nu-\beta_{1}x)y+z+G(x,y,z,\mu,\nu),$
$z^{\prime}$ $=(\nu-\beta_{1}x)z-y+H(x,y,z,\mu,\nu),$
for some new real-analytic functions $F$, $G$ and $H$ and new parameters $\beta_{1}$ and $\gamma_{2}$. Finally, we notice that the auxiliary equation
$x^{\prime}$ $=x^{2}-\mu+F(x,0,0,\mu,\nu),$
obtained by setting $y=z=0$ in the $x$-equation, is an equation for an analytic unfolding of a planar saddle-node. Here it is well-known that the $x^{3}$-term of $F(x,0,0,0,0)$ cannot be removed by normal form transformations, see e.g. [44]. We consider the expansion
$F(x,0,0,\mu,\nu)=\sum_{\alpha+\beta+\gamma=3}F_{\alpha,\beta,\gamma}x^{\alpha}\mu^{\beta}\nu^{\gamma}+\mathcal{O}(|(x,\mu,\nu)|^{4}),$
with $\mathcal{O}(|(x,\mu,\nu)|^{4})$ denoting terms that are fourth order with respect to $(x,\mu,\nu)\rightarrow(0,0,0)$. It is then a simple calculation to show that
$x=\widetilde{x}-\frac{1}{2}F_{1,2,0}\mu^{2}-\frac{1}{2}F_{1,0,2}\nu^{2}-\frac{1}{2}F_{1,1,1}\mu\nu-F_{2,1,0}\widetilde{x}\mu-F_{2,0,1}\widetilde{x}\nu$ (A.3)
which defines a local $(\mu,\nu)$-dependent change of coordinates $\widetilde{x}\mapsto x$, brings the auxiliary equation into the following form
$\widetilde{x}^{\prime}=\widetilde{x}^{2}-\widetilde{\mu}(\mu,\nu)+F_{3,0,0}x^{3}+\mathcal{O}(|(x,\mu,\nu)|^{4}),$
with $\widetilde{\mu}(\mu,\nu)=\mu+\mathcal{O}(2)$ an analytic function. Therefore by applying the $(\mu,\nu)$-dependent change of coordinates $\widetilde{x}\mapsto x$ defined by (A.3) to the full system (A.2), and a subsequent diffeomorphic change of parameters $(\mu,\nu)\mapsto(\widetilde{\mu},\nu)$, we again obtain (A.2) upon dropping the tildes, but where the new $F$ now satisfies
$F(x,0,0,\mu,\nu)=F_{3,0,0}x^{3}+\mathcal{O}(|(x,\mu,\nu)|^{4}).$ (A.4)
To complete the derivation of (1.1) satisfying (1.3), we simply set
$\begin{cases}\gamma_{2}=a,\\ \beta_{1}=b,\\ F_{3,0,0}=F_{0}.\end{cases}$
Remark A.1.
It is an interesting problem whether $\mathcal{O}(|(x,\mu,\nu)|^{4})$ in (A.4) can be removed completely within the analytic case. The recent paper [23] studies a similar problem in the smooth topology and seem to obtain partial results of this kind.